### v2

# Large-$N$ SU(4) Schwinger boson theory for coupled-dimer antiferromagnets

Abstract
We develop a systematic large-$N$ expansion based on the Schwinger boson representation of SU(4) coherent states of dimers for the paradigmatic spin-$1/2$ bilayer square lattice Heisenberg antiferromagnet. This system exhibits a quantum phase transition between a quantum paramagnetic state and a Néel order state, driven by the coupling constant $g=J^{\prime}/J$, which is defined as the ratio between the inter-dimer $J^{\prime}$ and intra-dimer $J$ exchange interactions. We demonstrate that this approach accurately describes static and dynamic properties on both sides of the quantum phase transition. The critical coupling constant $g_{c}\approx 0.42$ and the dynamic spin structure factor reproduce quantum Monte Carlo results with high precision. Notably, the $1/N$ corrections reveal the longitudinal mode of the magnetically ordered phase along with the overdamping caused by its decay into the two-magnon continuum. The present large-$N$ $SU(N)$ Schwinger boson theory can be extended to more general cases of quantum paramagnets that undergo a quantum phase transition into magnetically ordered states.
I Introduction
Networks of antiferromagnetic (AFM) dimers provide the simplest realization of quantum phase transitions between a quantum paramagnet (QPM) phase, adiabatically connected to a direct product of singlet states in each dimer, and a magnetically ordered phase [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The elementary excitations of the QPM states are spin-one quasi-particles known as triplons. These arise from exciting one singlet dimer into a triplet state that propagates in a periodic dimer lattice with well-defined momentum. When the inter-dimer interaction $J^{\prime}$ is much weaker than the intra-dimer exchange $J$, the single-triplon dispersion exhibits a spin gap of the order of $J$. As the ratio $g=J^{\prime}/J$ increases, the gap reduces and often vanishes continuously at a critical value $g=g_{c}$, signaling the onset of long-range magnetic ordering via condensation of the soft triplon mode [11, 12]. The magnetic ordering wave vector coincides with the wave vector of the triplon mode that becomes soft. While this simple and intuitive picture has been validated by multiple analytical, numerical, and experimental studies of quantum dimer magnets, an accurate and controlled theoretical description of these magnets has yet to be achieved.
A paradigmatic example is a spin-$1/2$ bilayer square lattice Heisenberg antiferromagnet [13], where $J^{\prime}$ connects nearest-neighbor spins on the same layer. This unfrustrated quantum model is free from the sign problem, allowing for accurate quantum Monte Carlo (QMC) simulations [14, 15, 16]. QMC studies reveal a second order quantum phase transition (QPT) between a QPM and a Néel ordered AFM phase at $g_{c}\approx 0.3965$. The QPT belongs to the 3D $O(3)$ universality class, where the spin gap closes as $\Delta=|g-g_{c}|^{\nu}$, with a critical exponent $\nu\approx 0.71$ [15].
Several other approaches have been employed to solve the bilayer antiferromagnet  [17, 18, 19, 20, 21, 22, 23, 24, 25, 16, 26, 27, 28]. From the magnetically ordered side, conventional linear spin-wave theory and its modified self-consistent versions, along with the Schwinger boson (SB) $SU(2)$ mean-field approximation, have incorrectly predicted a weak first-order transition at much smaller critical values of $g_{c}$ (see Table  1), in stark contrast with QMC results. Chubukov and Morr [17] identified that the failure of these theories was due to the neglect of longitudinal fluctuations, specifically those along the local magnetization, which are expected to be significantly enhanced near the QPT.
Approaching the problem from the quantum paramagnetic side, Chubukov [29], and later Sachdev and Bhatt [30], introduced bond-operator theories (BOT), where spin operators within each dimer are expressed as bilinear forms of boson operators. These bond operators correspond to the four levels of a dimer: one singlet and three triplets. The BOT approach relies on correctly accounting for the degrees of freedom within each dimer, naturally incorporating intra-dimer entanglement, and has been applied to both the QPM and AFM phases. However, the BOT has a significant limitation due to the requirement of a local constraint on the number of bosons.
Broadly, three alternatives have been proposed to handle the constraint: (a) imposing it on average with a Lagrange multiplier within a mean-field approach where the singlet operator is replaced by a $c$-number [14, 20]; (b) projecting out the double occupancy of triplets in each dimer [19, 31, 32]; and (c) using the constraint to express the singlet operator in terms of the triplet operators, akin to a Holstein-Primakoff transformation [33].
The mean-field treatment [20] -case (a)— predicts a critical value $g_{c}$ more in line with the QMC result (see Table 1). However, a significant shortcoming of this approach is that the solution of the mean-field equations yields a spectrum for the Néel phase that violates the Goldstone theorem [34]. We attribute these problems to the absence of an expansion parameter.
This limitation of the bond-operator theory was addressed by introducing a systematic perturbation scheme [31, 32]-case (b)-, by reformulating the bond-operator approach as an expansion in $1/d$, where $d$ is the spatial dimension of the system. While this method preserves the Goldstone modes of the AFM phase order by order in $1/d$ and also predicts a continuous phase transition, the resulting value of $g_{c}$ is no longer close to the QMC result (see Table 1).
The four bond operators can be identified with the four Schwinger bosons associated with the fundamental irreducible representation (irrep) of SU(4) [30], as their bilinears faithfully represent the $\mathrm{SU}(4)$ generators of infinitesimal unitary transformations within the local four-dimensional Hilbert space of each dimer. This representation can be extended to any completely symmetric irrep of $\mathrm{SU}(4)$ by adjusting the local constraint from one to $M$ bosons per dimer [35], where $M$ is a positive integer that labels different irreps. Notably, the classical limit of this theory is obtained by taking the large $M$ limit [36]. In this context, $M$ plays a role analogous to the spin value $S$ in the conventional SU(2) case.
A semi-classical (loop) expansion can be developed by implementing a Holstein-Primakoff transformation and expanding the bosonic propagators in powers of $1/M$, namely a generalized spin wave theory [37, 36] -case (c)-. However, to leading order, this large-$M$ expansion predicts a critical value $g_{c}\simeq 0.25$ which significantly deviates from the QMC result for the bilayer square lattice antiferromagnet.
The SU(2) Schwinger boson theory can be improved by applying a Gutzwiller projection [38] to enforce the local constraint. This procedure results in a better critical value of $g_{c}$ than the mean-field result, although $g_{c}$ remains significantly smaller than the QMC result (see Table 1). Another possibility is to perform a $1/N$-correction of the mean-field (saddle-point) solution. However, both, the Gutzwiller projection and large-$N$ expansion, start from an SU(2) Schwinger boson representation of the spin $S=1/2$ operators.
Method      $g_{c}\;\;\;$ Transition Goldstone Ref
QMC 0.3965 2nd yes [16]
LSWT 0.0735 1st yes [17]
MSWT 0.236 1st yes [39]
SBMFT 0.2232 1st yes [40]
SB Gutzwiller 0.285 2nd - [38]
BOT - $\lambda$ 0.437 2nd no [20]
BOT - Brueckner 0.389 - - [19]
large-$d$ BOT 0.2968 2nd yes [32]
SE 0.3942 2nd yes [18]
CST 0.382 2nd - [28]
large-$M$ SU(4) HP 0.25 2nd yes [16]
large-$N$ SU(4) SB 0.42 2nd yes This work
Table 1: Representative numerical and analytical results for the bilayer Heisenberg quantum antiferromagnet. $g_{c}$ indicates the critical coupling, “Transition” the order of the phase transition, "Goldstone“ indicates if the theory captures or not the Goldstone modes of the magnetically ordered phase. QMC stands for Quantum Monte Carlo, LSWT for linear spin wave theory, MSWT for modified spin wave theory, SBMFT for SU(2) Schwinger boson mean-field theory, SB Gutzwiller for SBMFT Gutzwiller projected, BOT-$\lambda$ for bond operator mean-field theory with the constraint imposed by a Lagrange multiplier, BOT+Brueckner for bond-operator theory with a constraint of no triplet double-occupancy, large-$d$ BOT for the $1/d$ expansion of the BOT, SE for dimer series expansions, CST for continuous similarity transformations, large-$M$ SU(4) HP for the generalized Holstein-Primakoff spin wave theory (equivalent to large-$d$ BOT for $d=\infty$), and large-$N$ SU(4) SB corresponds to this work.
We emphasize the asymmetry between the aforementioned semi-classical large-$M$ approach and the above-mentioned large-$N$ method. While the former is implemented using SU(4) coherent states that capture intra-dimer entanglement at the classical level, the large-$N$ methods that have been implemented so far are based on SU(2) coherent states for each site of the dimer, requiring intra-dimer entanglement to be incorporated via fluctuations beyond the saddle-point (SP). One way to address this asymmetry is to implement the latter using the same SU(4) Schwinger bosons that form the basis of the large-$M$ expansion. By capturing intra-dimer entanglement at the SP level of the SU(4) Schwinger boson theory, we can expect better agreement with numerical results.
In this work, we develop a systematic large-$N$ expansion based on the $\mathrm{SU}(4)$ coherent state description of a single dimer. This approach generalizes the original dimer problem with two antiferromagnetically coupled $\mathrm{SU}(2)$ spins on each site of the dimer to two antiferromagnetically coupled $\mathrm{SU}(n)$ spins, with $n\geq 2$. Unlike the traditional SB approximation, which usually employs $\mathrm{SU}(n)$ Schwinger bosons to represent the $\mathrm{SU}(n)$ spin components [41], we start from $\mathrm{SU}(N=n^{2})$ SBs to include all quantum mechanical states of the generalized dimer in the manifold of $\mathrm{SU}(N)$ coherent states. Our theory gives rise to a controlled structure and a $1/N$ expansion of the BOT mean-field approximation, where the boson number constraint is imposed by a Lagrange multiplier.
We compute the dynamical susceptibility and take the limit $N=n^{2}\to\infty$, using $1/N$ as the small expansion parameter. Remarkably, the leading-order SP contribution already provides a very accurate description of both phases, QPM and AFM. The low energy spectrum of the QPM consist of a triplet of gapped triplon modes that become soft at the critical point, $g=g_{c}$. The triplon condensation leads to the AFM state for $g>g_{c}$, where the transverse component of the dynamical spin structure factor (DSSF) is also accurately described by the SP approximation. Importantly, the DSSF exhibits magnon excitations with the expected transverse Goldstone modes and the value of $g_{c}\approx 0.42$ aligns well with QMC simulations. Moreover, the excitation spectrum and the spectral weight distribution of the dynamic spin structure factor (DSSF) also reproduce the QMC results with high precision.
Notably, the longitudinal mode of the magnetically ordered phase is also captured by the large-$N$ approach by including $1/N$ corrections beyond the SP level. These corrections account for the overdamping caused by the decay of the longitudinal mode into the two-magnon continuum. These results, which also align well with existing numerical simulations, indicate that the large-$N$ approach effectively captures the essential physics of this system.
This paper is organized as follows. In Sec. II, we introduce the $\mathrm{SU}(4)$ coherent states of a single dimer. Sec. III presents the Schwinger boson theory for coupled dimer antiferromagnets based on these $\mathrm{SU}(4)$ coherent states. We generalize the theory by extending the $\mathrm{SU}(4)$ formulation to $\mathrm{SU}(N=n^{2})$, corresponding to models of two antiferromagnetically coupled $\mathrm{SU}(n)$ spins on each dimer. A systematic expansion in powers of $1/N$ is then discussed within the path-integral formulation. In Sec. IV, we explore the SP solution of the large-$N$ theory, presenting the continuous transition between the QPM and the Néel AFM order. In Sec. V, we study the DSSF using the large-$N$ framework. Notably, we find that the SP approximation of the DSSF in the QPM phase and the transverse DSSF in the AFM phase agree very well with the QMC results (Sec. V.1). A correct description of the longitudinal DSSF requires the consideration of $1/N$ diagrams, which we discuss in Sec. V.2. Finally, the main conclusions of this study are summarized in Sec. VI.
II $\mathrm{SU}(4)$ coherent-state description of a dimer
We consider a system of antiferromagnetically coupled dimers. Each dimer consists of two $S=1/2$ spins coupled through an antiferromagnetic Heisenberg interaction, described by the following Hamiltonian
^$\hat{\mathscr{H}}_{0}=J\sum_{j}\left(\hat{\bm{S}}_{j+}\cdot\hat{\bm{S}}_{j-}-\frac{1}{4}\right)$ (1)
The index $j$ labels the dimers, while $\pm$ refers to the two spins of each dimer (see Fig. 1(a)). The eigenstates of a single dimer consist of a singlet ground state and three degenerate triplet states, whose wave functions and energy eigenvalues are
$|S=0\rangle_{j}$ $=$ $\frac{1}{\sqrt{2}}(|\uparrow\downarrow\rangle-|\downarrow\uparrow\rangle),\quad\quad\epsilon_{s}=-J,$
$|S=1,S^{x}=0\rangle_{j}$ $=$ $\frac{1}{\sqrt{2}}(|\downarrow\downarrow\rangle-|\uparrow\uparrow\rangle),\quad\quad\epsilon_{t,x}=0,$
$|S=1,S^{y}=0\rangle_{j}$ $=$ $\frac{i}{\sqrt{2}}(|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle),\quad\quad\epsilon_{t,y}=0,$
$|S=1,S^{z}=0\rangle_{j}$ $=$ $\frac{1}{\sqrt{2}}(|\uparrow\downarrow\rangle+|\downarrow\uparrow\rangle),\quad\quad\epsilon_{t,z}=0.$ (2)
In the absence of inter-dimer interaction, the ground state of the system is a product state of singlets on each dimer, and the excitations are local flips from a singlet to triplet (i.e., triplon excitation), with an energy gap $J$. This solvable limit provides a qualitative picture for finite but weak inter-dimer interaction. The ground state is still a quantum paramagnet and the quantum mechanical state of each dimer retains a strong singlet character. In presence of translational invariance, the elementary excitations are triplon modes with well-defined momentum that propagate through the lattice.
The goal of this work is to understand the system’s behavior upon increasing the inter-dimer interaction. Since the system may develop different types of instabilities due to the softening of the triplon modes, it is important to introduce a formalism that can quantitatively describe these QPTs, as well as the low-energy excitation spectrum on both sides of the transition.
A crucial consideration in choosing an adequate formalism is to note that each dimer is an “entangled unit” in the sense that the corresponding wave function has a strong singlet character. It is then convenient to use a formalism in which the singlet state of each dimer is a coherent state of the Lie algebra associated with bosonic operators that are used to represent the spin operators. More specifically, a standard approach to the problem would start with a faithful representation of the spin operators at each site $j\sigma$ ($\sigma=\pm$ denotes the layer) in terms of $\mathrm{SU}(2)$ SB $\hat{b}^{\dagger}_{j\sigma,\mu}$ and $\hat{b}^{\;}_{j\sigma,\mu}$ ($\mu=\uparrow\downarrow$),
$\hat{S}^{+}_{j\sigma}$ $=$ $\hat{b}^{\dagger}_{j\sigma,\uparrow}\hat{b}^{\;}_{j\sigma,\downarrow},\quad\hat{S}^{-}_{j\sigma}=\hat{b}^{\dagger}_{j\sigma,\downarrow}\hat{b}^{\;}_{j\sigma,\uparrow},$
$\hat{S}^{z}_{j\sigma}$ $=$ $\frac{1}{2}(\hat{b}^{\dagger}_{j\sigma,\uparrow}\hat{b}^{\;}_{j\sigma,\uparrow}-\hat{b}^{\dagger}_{j\sigma,\downarrow}\hat{b}^{\;}_{j\sigma,\downarrow}),$ (3)
that fulfill the constraint:
$\hat{b}^{\dagger}_{j\sigma,\downarrow}\hat{b}^{\;}_{j\sigma,\downarrow}+\hat{b}^{\dagger}_{j\sigma,\uparrow}\hat{b}^{\;}_{j\sigma,\uparrow}=2S,$ (4)
where $S$ labels the irrep of $\mathrm{SU}(2)$ that determines the matrix form of the generators given in Eq. (3). Note that $S=1/2$ for the particular case of interest. However, as we mentioned in the Introduction, a large-$N$ approximation based on these SBs is not quantitatively accurate because the path-integral formulation is parametrized in terms of direct products of $\mathrm{SU}(2)$ coherent states. In other words, the intra-dimer entanglement must be built in via quantum fluctuations that favor linear combination of these product states.
The above-mentioned problem can be avoided by introducing bosons with four flavours (instead of two), $\hat{b}^{\dagger}_{j,\mu}$ and $\hat{b}^{\;}_{j,\mu}$ ($\mu=0,1,2,3$), that fulfill the constraint
$\sum_{\mu=0}^{3}\hat{b}^{\dagger}_{j,\mu}\hat{b}^{\;}_{j,\mu}=M,$ (5)
where $M$ is a positive integer ($M=1,2,3,...$). Bilinear forms in these bosons,
$\hat{\cal S}^{\mu\nu}_{j}=\hat{b}^{\dagger}_{j,\mu}\hat{b}_{j,\nu},$ (6)
with commutation relations
$[\hat{\cal S}^{\alpha\beta}_{j},\hat{\cal S}^{\mu\nu}_{j}]=\delta_{\beta\mu}\hat{\cal S}^{\alpha\nu}_{j}-\delta_{\alpha\nu}\hat{\cal S}^{\mu\beta}_{j},$ (7)
provide a faithful representation of generators of $\mathrm{SU}(4)$ in the completely symmetric irreps labelled by the integer $M$ (e.g. $M=1$ for the fundamental irrep of $\mathrm{SU}(4)$). Note that unlike the $\mathrm{SU}(2)$ SBs, these bosons have only a dimer index $j$ because they create all the possible quantum mechanical states of each dimer. In other words, the $\mathrm{SU}(4)$ SBs create coherent states of the $\mathrm{SU}(4)$ Lie algebra (for completely symmetric irreps) that span the CP3 manifold of quantum mechanical states of a four-level quantum mechanical system. To make contact with our problem of interest, we can regard the single dimer $j$ as the four-level system, set $M=1$ (fundamental irrep of $\mathrm{SU}(4)$) and choose the basis of $\mathrm{SU}(4)$ SBs, such that the $\mu=0$ boson creates the singlet state:
$|S=0\rangle_{j}=\hat{b}^{\dagger}_{j,0}|0\rangle,$ (8)
while the other three bosons $\mu=1,2,3$ create the three triplet states (corresponding to the index $\mu=x,y,z$ )
$|S=1,S^{\mu}=0\rangle_{j}=\hat{b}^{\dagger}_{j,\mu}|0\rangle.$ (9)
Figure 1: (a) A single dimer formed by two $S=1/2$ spins coupled through antiferromagnetic interaction $J$. (b) Network of dimers forming a square lattice. The splitting between sublattice ${\cal A}$ (white circles) and ${\cal B}$ (black circles) of the dimer lattice becomes relevant in presence of Néel AFM order.
The important difference between the $\mathrm{SU}(2)$ and the $\mathrm{SU}(4)$ SBs is that coherent states of the latter include intra-dimer singlet states, meaning that the intra-dimer entanglement is already built-in the $\mathrm{SU}(4)$ coherent states of the path-integral formulation and only the inter-dimer entanglement must be generated via the inclusion of quantum fluctuations. As expected from this simple observation, we will see in the next sections that the $\mathrm{SU}(4)$ SB theory (large-$N$ expansion) is indeed much more accurate than the $\mathrm{SU}(2)$ SB theory for low-dimensional coupled-dimer systems.
III $\mathrm{SU}(4)$ Schwinger boson theory
To be specific, we shall consider a double-layer square lattice antiferromagnets, where two $S=1/2$ spins on the vertical inter-layer bonds connecting two layers form a antiferromagnetic dimer (see Fig. 1 (b)). The spin Hamiltonian is the sum of the intra-dimer and inter-dimer spin Hamiltonians ^$\hat{\mathscr{H}}_{0}$ and $\hat{\mathscr{H}}^{\prime}$, respectively:
^=^0+^,^subscript^$\hat{\mathscr{H}}=\hat{\mathscr{H}}_{0}+\hat{\mathscr{H}}^{\prime},$ (10)
where
^$\hat{\mathscr{H}}^{\prime}=J^{\prime}\sum_{j,\delta}\sum_{\sigma=\pm}\hat{\bm{S}}_{j,\sigma}\cdot\hat{\bm{S}}_{j+\delta,\sigma},$ (11)
with $j$ being the dimer index, and $\delta=x,y$ running over the two inter-bond links associated with dimer $j$, and $j+\delta$ denoting the neighboring dimer connected to dimer $j$ through bond $\delta$. $J^{\prime}$ is the inter-dimer spin-exchange illustrated in Fig. 1(b). Up to the overall energy scale $J$, this problem is solely characterized by the dimensionless coupling constant $g=J^{\prime}/J$.
Since the $\mathrm{SU}(4)$ generators in the fundamental irrep ($M=1$), together with the identity, form a complete basis for the vector space of operators acting on the 4-dimensional Hilbert space of a single dimer, the spin operators on each site of the dimer can be expressed as a linear combination of these generators
$\hat{S}_{j\pm}^{\mu}$ $=$ ±$\pm\frac{1}{2}\left(\hat{\cal S}_{j}^{0\mu}+\hat{\cal S}_{j}^{\mu 0}\right)-\frac{i}{2}\sum_{\nu,\rho=1}^{3}\epsilon^{\mu\nu\rho}\hat{\cal S}_{j}^{\nu\rho}.$ (12)
In terms of the $\mathrm{SU}(4)$ SB, the single-dimer Hamiltonian takes the diagonal form:
^$\hat{\mathscr{H}}_{0}=-J\sum_{j}\hat{\cal S}^{00}_{j}=-J\sum_{j}\hat{b}^{\dagger}_{j,0}\hat{b}^{\;}_{j,0}.$ (13)
Given the local constraint in Eq. (5), the ground state $|\Psi_{0}\rangle$ of ^$\hat{\mathscr{H}}_{0}$ (direct product of singlet states) satisfies $\hat{b}_{j,0}^{\dagger}\hat{b}_{j,0}|\Psi_{0}\rangle=|\Psi_{0}\rangle$ for $\forall j$.
The inter-dimer interaction $\hat{{\mathscr{H}}}^{\prime}$ generates quantum fluctuations that convert a pair of singlet states into pairs of triplets that form a inter-dimer singlet. In addition, the low-energy excitations that were localized triplet modes in absence of $\hat{{\mathscr{H}}}^{\prime}$, become extended triplon modes because finite inter-dimer coupling can exchange a singlet and a triplet state in two different dimers. Both effects, creation of virtual triplon pairs and effective triplon hopping produce a finite dispersion of the still gapped single-triplon mode.
In the $\mathrm{SU}(4)$ SB theory, we rewrite the inter-dimer Hamiltonian in terms of $\mathrm{SU}(2)$-invariant “link” operators  [^1] In the literature, these are usually called “bond operators”. In this work we will call them “link operators” to distinguish them from the “bond operators” acting on each dimer unit.
$\hat{\mathscr{H}}^{\prime}$ $=$ $\frac{J^{\prime}}{2}\sum_{j,\delta}\left(\hat{S}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}+\hat{T}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}+\mathrm{h.c.}\right)$ (14)
$+$ $\frac{J^{\prime}}{2}\sum_{j,\delta}\left(:\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}:-\hat{A}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}\right),$
where $:\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}:$ indicates normal ordering of the operator $\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}$, and the four $\mathrm{SU}(2)$ invariant link operators are
$\hat{S}_{j,\delta}=\hat{b}_{j,0}\hat{b}_{j+\delta,0},\qquad\ \hat{T}_{j,\delta}^{\dagger}=\hat{b}_{j,0}^{\dagger}\hat{b}_{j+\delta,0},$ (15)
$\hat{A}_{j,\delta}=\sum_{\mu=1}^{3}\hat{b}_{j,\mu}\hat{b}_{j+\delta,\mu},\ \hat{B}_{j,\delta}^{\dagger}=\sum_{\mu=1}^{3}\hat{b}_{j,\mu}^{\dagger}\hat{b}_{j+\delta,\mu}.$ (16)
We note that the singlet boson $\hat{b}_{j,0}$ remains invariant (trivial irrep) under a global $\mathrm{SU}(2)$ rotation of the dimer $j$, while the triplet bosons $\hat{b}_{j,\mu}$ ($\mu=1,2,3$) transform according to the $L=1$ (adjoint) irrep. The $\mathrm{SU}(2)$ invariance of $\hat{S}_{j,\delta}$ and $\hat{T}_{j,\delta}^{\dagger}$ follows directly from the singlet character of $\hat{b}^{\dagger}_{j,0}$ and $\hat{b}_{j,0}$. The other link operators, $\hat{A}_{j,\delta}$ and $\hat{B}_{j,\delta}^{\dagger}$, are the two singlets obtained by considering the two possible direct products of bosons that transform like triplets, $1\otimes 1=0\oplus 1\oplus 2$, and projecting into the singlet $(0)$ component.
$\hat{\mathscr{H}}^{\prime}$ (14) is expressed in a form that is manifestly SU(2) symmetric. Physically, the different terms of $\hat{\mathscr{H}}^{\prime}$ can be interpreted in the following way: $\hat{A}^{\dagger}\hat{S}$ represents the process of destroying two neighboring singlets followed by the creation of two triplets, which together form a singlet (${\rm h.c.}$ corresponds to the inverse process); $\hat{T}^{\dagger}\hat{B}$ the exchange between singlet and triplet neighbours, $:\hat{B}^{\dagger}\hat{B}:$ the exchange between neighboring triplets, and $\hat{A}^{\dagger}\hat{A}$ projects onto singlet formed between two neighboring triplets.
For strong enough inter-dimer interaction, the system undergoes a transition to a Néel AFM state, where the magnetic moments on nearest-neighbors within the same layer or within the same dimer are anti-aligned. Since this magnetic order breaks translation symmetry, we switch to a twisted reference frame in which the ground state remains translationally invariant and the magnetic ordering is ferromagnetic in each layer, but antiferromagnetic between layers. To define the twisted reference frame it is convenient to introduce two interpenetrated sublattices ${\cal A}$ and ${\cal B}$ of dimers (see Fig. 1). The twisted reference frame is obtained by applying a $\pi$-rotation about the spin-$z$ axis on spins of the ${\cal B}$ sublattice: $S_{j\sigma}^{x/y}\to-S_{j\sigma}^{x/y}$, $S_{j\sigma}^{z}\to S_{j\sigma}^{z}$. Under this transformation, the singlet and the three triplet states in Eq. (2) transform in the following way:
$|S=0\rangle_{j}$$\to$ $|S=0\rangle_{j},$
$|S=1,S^{z}=0\rangle_{j}$$\to$ $|S=1,S^{z}=0\rangle_{j},$
$|S=1,S^{x}=0\rangle_{j}$$\to$ $e^{i\bm{\pi}\cdot\bm{r}_{j}}|S=1,S^{x}=0\rangle_{j},$
$|S=1,S^{y}=0\rangle_{j}$$\to$ $e^{i\bm{\pi}\cdot\bm{r}_{j}}|S=1,S^{y}=0\rangle_{j},$ (17)
[…]
$\hat{b}_{j,0}\rightarrow\hat{b}_{j,0},\quad\hat{b}_{j,1}$$\rightarrow$ $\hat{b}_{j,1}e^{i\bm{\pi}\cdot\bm{r}_{j}},$
$\hat{b}_{j,3}\rightarrow\hat{b}_{j,3},\quad\hat{b}_{j,2}$$\rightarrow$ $\hat{b}_{j,2}e^{i\bm{\pi}\cdot\bm{r}_{j}}.$ (18)
In the twisted reference frame, the SB representation of the spin operators is still given by Eq. (12), while the four link operators transform to
$\hat{S}_{j,\delta}$ $=\hat{b}_{j,0}\hat{b}_{j+\delta,0},\ \ \hat{T}_{j,\delta}^{\dagger}=\hat{b}_{j,0}^{\dagger}\hat{b}_{j+\delta,0},$ (19)
$\hat{A}_{j,\delta}$ $=\hat{b}_{j,3}\hat{b}_{j+\delta,3}-\hat{b}_{j,2}\hat{b}_{j+\delta,2}-\hat{b}_{j,1}\hat{b}_{j+\delta,1},$ (20)
$\hat{B}_{j,\delta}^{\dagger}$ $=\hat{b}_{j,3}^{\dagger}\hat{b}_{j+\delta,3}-\hat{b}_{j,2}^{\dagger}\hat{b}_{j+\delta,2}-\hat{b}_{j,1}^{\dagger}\hat{b}_{j+\delta,1}.$ (21)
Unless otherwise specified, we will operate within this twisted reference frame.
III.1 Large-$N$ generalization
To study this interacting boson system, we use the large-$N$ technique, which extends the original $\mathrm{SU}(2)$ invariant system to a system with $\mathrm{SU}(n)$ symmetry. Each dimer consists of two antiferromagnetically coupled $\mathrm{SU}(n)$ spins described by:
^$\hat{\mathscr{H}}_{0}=\frac{4\tilde{J}(N)}{N}\sum_{j}\left(\hat{O}_{j,+}^{\mu}\hat{\tilde{O}}_{j,-}^{\mu}+C(N)\right),$ (22)
where $N=n^{2}$, $\hat{O}^{\mu}_{j,+}$ denotes the generators of the $\mathrm{SU}(n)$ group at the “$+$” site of the dimer in the fundamental irrep, while $\hat{\tilde{O}}_{j,-}^{\mu}$ are the $\mathrm{SU}(n)$ generators in the conjugate representation for the “--” site. $C(N)$ is an irrelevant constant that is used to set the highest energy eigenvalue to zero. The tensor product of these representations is reducible, i.e., ${n}\otimes\bar{{n}}={1}\oplus{n^{2}-1}$, where the numbers refer to the dimensions of the corresponding irrep. We associate the $N$ flavors of the bosons with the energy levels of each dimer, denoted as $\hat{b}_{j,\sigma}$, where $\sigma=0,1,2,...,N-1$. Here, $\hat{b}_{j,0}^{\dagger}$ creates an $\mathrm{SU}(n)$ singlet state with energy $-J$ that scales as $N^{0}$ in the large-$N$ limit. The remaining bosons generate an $N-1$ dimensional multiplet with zero energy, assuming appropriate choices of the generalized exchange constant $\tilde{J}(N)$ and the global shift $C(N)$. The $\mathrm{SU}(n)$ generators are given by:
$\hat{O}^{\mu}_{j,+}=\frac{1}{\sqrt{2n}}(\hat{b}_{j,0}^{\dagger}\hat{b}_{j,\mu}+\mathrm{h.c.})+\sum_{\nu,\rho=1}^{N-1}\frac{d_{\mu\nu\rho}-if_{\mu\nu\rho}}{2}\hat{b}_{j,\nu}^{\dagger}\hat{b}_{j,\rho},$ (23)
and
$\hat{\tilde{O}}^{\mu}_{j,-}=\frac{-1}{\sqrt{2n}}(\hat{b}_{j,0}^{\dagger}\hat{b}_{j,\mu}+\mathrm{h.c.})-\sum_{\nu,\rho=1}^{N-1}\frac{d_{\mu\nu\rho}+if_{\mu\nu\rho}}{2}\hat{b}_{j,\nu}^{\dagger}\hat{b}_{j,\rho},$ (24)
where $11\leq\mu\leq N-$ 1 $,f_{\mu\nu\rho}$ are the structure constants of the $\mathfrak{su}(n)$ Lie algebra (completely antisymmetric under exchange of any pair of indices), and $d_{\mu\nu\rho}$ is a fully symmetric third rank tensor, defined by the anticommutator of the $\mathrm{SU}(n)$ generators:
$d_{\mu\nu\rho}=2\operatorname{Tr}\left[\left\{T^{\mu},T^{\nu}\right\}T^{\rho}\right],$ (25)
where the index $11\leq\mu,\nu,\rho\leq N-$ 1. For the $\mathrm{SU}(2)$ case, $d_{\mu\nu\rho}\equiv 0$ and $f_{\mu\nu\rho}=\epsilon_{\mu\nu\rho}$, so we recover the expression for spin operators in Eq. (12).
In this generalization of the model to arbitrary $n$ values, with $N=n^{2}$, the isolated dimer Hamiltonian retains the form of the original spin-$1/2$ Hamiltonian, as defined in Eq. (13). The extension of the inter-dimer interaction is based on Eq. (14), namely,
$\hat{\mathscr{H}}^{\prime}$ $=$ $\frac{2J^{\prime}}{N}\sum_{j,\delta}\left(\hat{S}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}+\hat{T}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}+\mathrm{h.c.}\right)$ (26)
$+\frac{2J^{\prime}}{N}\sum_{j,\delta}\left(:\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}:-\hat{A}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}\right),$
where the link operators
$\hat{S}_{j,\delta}=\hat{b}_{j,0}\hat{b}_{j+\delta,0},\quad\hat{T}_{j,\delta}=\hat{b}_{j,0}^{\dagger}\hat{b}_{j+\delta,0},$ (27)
are bilinears in singlet bosons, while the link operators
$\hat{A}_{j,\delta}=\sum_{\mu=1}^{N-1}\hat{b}_{j,\mu}\hat{b}_{j+\delta,\mu},\quad\hat{B}_{j,\delta}=\sum_{\mu=1}^{N-1}\hat{b}_{j,\mu}^{\dagger}\hat{b}_{j+\delta,\mu},$ (28)
consist of bilinears in the $N-1$ flavor bosons that transform according to the (self-conjugate) adjoint irrep of $\mathrm{SU}(N)$. Notably, all four bond operators are $\mathrm{SU}(n)$ invariant.
Lastly, the local constraint Eq. (5) is generalized to:
$\sum_{\mu=0}^{N-1}\hat{b}^{\dagger}_{j,\mu}\hat{b}_{j,\mu}=M,$ (29)
where the integer $M$ labels different completely symmetric irreps of the $\mathrm{SU}(N)$ group. The $M\to\infty$ limit corresponds to a classical limit based on $\mathrm{SU}(N)$ coherent states of the extended model and quantum corrections can be incorporated via a $1/M$ expansion [36, 43, 44]. In the large-$N$ approach discussed here, $M$ must be proportional to $N$ and the ratio $\kappa=M/N$ remains a finite number in the large-$N$ limit (see Fig. 2). In this large-$N$ limit, the average boson density for each flavor remains of order one. For our original problem, we have $N=4$ and $M=1$.
Figure 2: Large-$N$ versus large-$M$ limit of the extended model. Large-$M$ limit corresponds to $M\to\infty$ while fixing $N$ ($N=4$ for the case of interest). The large-$N$ (SB theory) corresponds to $N\to\infty$ under a fix ratio $\kappa=M/N$, i.e., along the dotted line.
III.2 Path-integral formulation
The path-integral formulation is implemented by introducing identities expressed in terms of the $\mathrm{SU}(4)$ coherent states of each dimer, rather than the more conventional approach using $\mathrm{SU}(2)$ coherent states of each spin [41, 45]:
${\cal Z}$ $=$ $\int{\cal D}[\bar{b}b\lambda]\exp\left[-\int_{0}^{\beta}d\tau\left(\sum_{j\mu}\bar{b}_{j,\mu}\partial_{\tau}b_{j,\mu}+\mathscr{H}(\bar{b},b)\right)\right.$ (30)
$\left.+i\sum_{j}\lambda_{j}(\sum_{\mu}\bar{b}_{j,\mu}b_{j,\mu}-\kappa N)\right],$
where $\tau$ denotes the imaginary time, $\beta$ the inverse of temperature, $\mathscr{H}(\bar{b},b)=\mathscr{H}_{0}(\bar{b},b)+\mathscr{H}^{\prime}(\bar{b},b)$ the Hamiltonian density, obtained from ^0+^subscript^$\hat{\mathscr{H}}_{0}+\hat{\mathscr{H}}^{\prime}$ by replacing the operators $\hat{b}_{j,\mu}^{\dagger},\hat{b}_{j,\mu}$ with complex fields $\bar{b}_{j,\mu},b_{j,\mu}$ that parametrize the CP3 manifold of $\mathrm{SU}(4)$ coherent states. The additional fields are the Lagrange multipliers $\lambda_{j}$, which enforce the local constraint given in Eq. (29).
To decouple the terms that are quartic in the SBs, we first rewrite the link-operator factorization of ${\mathscr{H}}^{\prime}$ (see Eq. (26)) in the matrix form
$\mathscr{H}^{\prime}={4\over N}\sum_{j,\delta}\Psi_{j,\delta}^{\dagger}{\cal J}_{j,\delta}\Psi_{j,\delta},$ (31)
where $\Psi_{j,\delta}\equiv(S_{j,\delta},T_{j,\delta},A_{j,\delta},B_{j,\delta})^{T}$ denotes a 4-component link field  [^2] These link-fields are usually called “bond-fields” in the literature. Here we use the name “link-fields” to distinguish them from the intra-dimer “bond-fields”., $\dagger$” denotes the conjugate transpose of a complex vector, and
${\cal J}_{j,\delta}={J^{\prime}\over 2}\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\\ 1&0&-1&0\\ 0&1&0&1\end{array}\right),$ (36)
is a positive-defined matrix. Next, a Hubbard-Stratonovich (HS) transformation is implemented by introducing the auxiliary complex field $W_{j,\delta}=(W_{j,\delta}^{(S)},W_{j,\delta}^{(T)},W_{j,\delta}^{(A)},W_{j,\delta}^{(B)})^{T}$, that are conjugate of the $\Psi$-fields  [47]:
${\cal Z}=\int{\cal D}[\bar{W}W\bar{b}b\lambda]\exp\left[-S(\bar{W},W,\bar{b},b,\lambda)\right],$ (37)
where the action reads
$S(\bar{W},W,\bar{b},b,\lambda)$ $=$ $\int_{0}^{\beta}d\tau\left[\left(\sum_{j\mu}\bar{b}_{j,\mu}\partial_{\tau}b_{j,\mu}+{\mathscr{H}}_{0}(\bar{b},b)\right)+\sum_{j,\delta}\left({N\over 4}\bar{W}_{j,\delta}{\cal J}_{j,\delta}W_{j,\delta}-\bar{W}_{j,\delta}{\cal J}_{j,\delta}\Psi_{j,\delta}+\Psi_{j,\delta}^{\dagger}{\cal J}_{j,\delta}W_{j,\delta}\right)\right.$ (38)
$+$ $\left.i\sum_{j}\lambda_{j}(\sum_{\mu}\bar{b}_{j,\mu}b_{j,\mu}-\kappa N)\right],$
and $\bar{W}_{j,\delta}$ refers to the conjugate of the link field $W_{j,\delta}$. The phase fluctuations of the auxiliary fields represent the emergent gauge fluctuations of the SB theory [41, 45, 47].
The new expression for the action is quadratic in the bosonic fields. In addition to the quadratic terms arising from ${\mathscr{H}}_{0}$ and from the on-site Lagrange multiplier $\lambda_{j}$, there are complex hopping/pairing link terms arising from inter-dimer interactions. The large-$N$ expansion is obtained after integrating out the bosonic variables,
$e^{-{N\over 4}S_{\rm eff}(\bar{W},W,\lambda)}=\int{\cal D}[\bar{b}b]e^{-S(\bar{W},W,\bar{b},b,\lambda)},$ (39)
and expanding the resulting effective action $S_{\rm eff}(\bar{W},W,\lambda)$ in the auxiliary fields $\bar{W}$, $W$, and $\lambda$ around the SP solution. The SP condition,
$\left.\frac{\delta S_{\rm eff}}{\delta W_{j,\delta}}\right\rvert_{\rm sp}=0,\ \ \left.\frac{\delta S_{\rm eff}}{\delta\bar{W}_{j,\delta}}\right\rvert_{\rm sp}=0,\ \ \left.\frac{\delta S_{\rm eff}}{\delta\lambda_{j}}\right\rvert_{\rm sp}=0,$ (40)
gives rise to the set of self-consistent equations
$W_{j,\delta}\rvert_{\rm sp}$ $=$ ${4\over N}\langle{\Psi}_{j,\delta}\rangle_{\rm sp},$
$\bar{W}_{j,\delta}\rvert_{\rm sp}$ $=$ $-{4\over N}\langle{\Psi}_{j,\delta}^{\dagger}\rangle_{\rm sp},$
$\kappa N$ $=$ $\sum_{\mu}\langle\bar{b}_{j,\mu}b_{j,\mu}\rangle_{\rm sp},$ (41)
where the average is taken using the SP action $S_{\rm sp}(\bar{b},b)\equiv S(\bar{W}_{\rm sp},W_{\rm sp},\bar{b},b,\lambda_{\rm sp})$:
$\langle O\rangle_{\rm sp}\equiv{\int{\cal D}[\bar{b}b]Oe^{-S_{\rm sp}}\over\int{\cal D}[\bar{b}b]e^{-S_{\rm sp}}}.$ (42)
Since the SP bosonic Hamiltonian must be hermitian, the SP solution must satisfy $\bar{W}_{j,\delta}\rvert_{\rm sp}=-(W_{j,\delta}\rvert_{\rm sp})^{*}$ (here * refers to complex conjugate) and $\lambda_{j}\rvert_{\rm sp}=-i\tilde{\lambda}_{j}\rvert_{\rm sp}$ with $\tilde{\lambda}_{j}\rvert_{\rm sp}\in\mathbb{R}$ [47]. These SP equations correspond to the self-consistent equations of the mean-field SB theory in the canonical formalism. One can verify that the solutions $W_{j,\delta}\rvert_{\rm sp}$, $\bar{W}_{j,\delta}\rvert_{\rm sp}$, and $\tilde{\lambda}_{j}\rvert_{\rm sp}$ scale as $N^{0}$, and consistently the energy spectrum of the bosons scales as $N^{0}$ as well.
To account for the fluctuations of the auxiliary link fields, we introduce
$\Phi_{j}\equiv\left(W_{j}^{T}-(W_{j}\rvert_{\rm sp})^{T},\bar{W}_{j}^{T}-(\bar{W}_{j}\rvert_{\rm sp})^{T},\lambda_{j}-\lambda_{j}\rvert_{\rm sp}\right)^{T},$ (43)
where $W_{j}\equiv((W_{j,x})^{T},(W_{j,y})^{T})^{T}$ and $\Phi_{j}^{\lambda}=\lambda_{j}-\lambda_{j}\rvert_{\rm sp}$ are the real fluctuations of the Lagrange multiplier fields. We denote the conjugate of $\Phi_{j}$ as
$\bar{\Phi}_{j}\equiv\left(\bar{W}_{j}^{T}-(\bar{W}_{j}\rvert_{\rm sp})^{T},W_{j}^{T}-(W_{j}\rvert_{\rm sp})^{T},\lambda_{j}-\lambda_{j}\rvert_{\rm sp}\right).$ (44)
For convenience, we also use the link variables $\Phi_{j,\delta}\equiv(W_{j,\delta}^{T}-(W_{j,\delta}\rvert_{\rm sp})^{T})^{T}$ and $\bar{\Phi}_{j,\delta}\equiv(\bar{W}_{j,\delta}^{T}-(\bar{W}_{j,\delta}\rvert_{\rm sp})^{T})$ to denote the fluctuation fields on bond $\delta$. Expanding the action around the SP yields:
$S(\bar{W},W,\bar{b},b,\lambda)$ $=$ $S_{\rm cl}+\int_{0}^{\beta}d\tau\left[\left(\sum_{j\mu}\bar{b}_{j,\mu}\partial_{\tau}b_{j,\mu}+{\mathscr{H}}_{\rm sp}(\bar{b},b)\right)+\sum_{j,\delta}\left({N\over 4}\bar{\Phi}_{j,\delta}{\cal J}_{j,\delta}\Phi_{j,\delta}-\bar{\Phi}_{j,\delta}{\cal J}_{j,\delta}\Psi_{j,\delta}+\Psi_{j,\delta}^{\dagger}{\cal J}_{j,\delta}\Phi_{j,\delta}\right)\right.$ (45)
$+$ $\left.i\sum_{j}\Phi_{j}^{\lambda}\sum_{\mu}\bar{b}_{j,\mu}b_{j,\mu}\right],$
where
$S_{\rm cl}=\frac{N\beta}{4}\left[\sum_{j,\delta}(W_{j,\delta}\rvert_{\rm sp})^{\dagger}{\cal J}_{j,\delta}W_{j,\delta}\rvert_{\rm sp}+4i\kappa\sum_{j}\lambda_{j}\rvert_{\rm sp}\right]$
refers to the “classical” interaction energy for the auxiliary fields, which is proportional to $N$, ${\mathscr{H}}_{\rm sp}(\bar{b},b)$ is the SP Hamiltonian density, which is also of order $N$:
${{\mathscr{H}}}_{\rm sp}$ $=$ ${{\mathscr{H}}}_{0}+\sum_{j,\delta}(W_{j,\delta}\rvert_{\rm sp})^{\dagger}{\cal J}_{j,\delta}{\Psi}_{j,\delta}+{\Psi}_{j,\delta}^{\dagger}{\cal J}_{j,\delta}W_{j,\delta}\rvert_{\rm sp}$ (46)
$+$ $\sum_{j}\tilde{\lambda}_{j}\rvert_{\rm sp}\sum_{\mu}\bar{b}_{j,\mu}{b}_{j,\mu}.$
The remaining components in $S(\bar{W},W,\bar{b},b,\lambda)$ represent the fluctuations of the auxiliary fields and their interaction with the bosons.
Figure 3: Building blocks for diagrammatic representation of the large-$N$ expansion. (a) Single-boson propagator at the SP level. (b) Internal vertex defined in Sec. III.2. (c) External vertex that couples the external sources to the bosonic propagator. (d) Polarization bubble and corresponding RPA propagator of the auxiliary fields.
Since we are working in the twisted reference frame, the expected SP solution is translationally invariant and the quadratic action becomes diagonal in momentum space. It is then convenient to rewrite the action in momentum space:
$S=S_{\rm cl}+S_{0}(\bar{\Phi},\Phi)+S_{1}(\bar{\Phi},\Phi,\bar{\psi},\psi),$ (47)
where $S_{\rm cl}$ has been defined above. The second term reads
$S_{0}={N\over 8}\sum_{q,\alpha,\alpha^{\prime}}(\bar{\Phi})_{q}^{\alpha}{\Pi}_{0}^{\alpha\alpha^{\prime}}\Phi_{q}^{\alpha^{\prime}},$ (48)
where $q\equiv(\bm{q},i\omega_{n})$, $\omega_{n}=2\pi n/\beta$ ($n\in{\mathbb{Z}}$) are the bosonic Matsubara frequencies,
[…]
the Fourier transform of the fluctuation fields, with ${\cal N}_{D}$ being the total number of dimers, and the block-diagonal constant matrix reads
${\Pi}_{0}=\left(\begin{array}[]{c|c|c|c|c}{\cal J}_{x}&&&&\\ \hline\cr&{\cal J}_{y}&&&\\ \hline\cr&&{\cal J}_{x}^{T}&&\\ \hline\cr&&&{\cal J}_{y}^{T}&\\ \hline\cr&&&&0\\ \end{array}\right),$ (56)
where ${\cal J}_{\delta}\equiv{\cal J}_{j,\delta}$ (independent to $j$) due to the translational invariance. The third term in $S$ reads
$S_{1}={1\over 2\sqrt{{\cal N}_{D}\beta}}\sum_{kq}\psi_{k}^{\dagger}\left(\delta_{q,0}G_{\rm sp}^{-1}(k)+2\sum_{\alpha}\Phi^{\alpha}_{-q}V^{\alpha}_{k,k+q}\right)\psi_{k+q}.$ (57)
where $k\equiv(\bm{k},i\nu_{m})$ $\nu_{m}=2\pi m/\beta$ ($m\in{\mathbb{Z}}$) is the bosonic Matsubara frequency, $G_{\rm sp}^{-1}(k)$ is the inverse of the Green’s function of the SB at the SP approximation,
$G_{\rm sp}^{-1}(k)=-i\nu_{m}\sigma_{z}\otimes I_{N}+H_{\rm sp}({\bm{k}}).$ (58)
with $I_{N}$ denoting the $N$-dimensional identity matrix and $H_{\rm sp}({\bm{k}})$ the SP Hamiltonian in momentum space, $\psi_{k}$ is the Nambu representation for the SB, defined by
$\psi_{j}$ $=$ $(b_{j}^{T},\bar{b}_{j})^{T},\quad b_{j}=(b_{j,0},...,b_{j,N-1})^{T},$ (59)
[…]
and $V^{\alpha}_{k+q,k}$ is the “internal” interaction vertex between the fluctuation fields and the boson field, whose diagrammatic representation is illustrated in Fig 3 (b). Explicitly, by defining the bilinear form of the link fields under the Nambu basis of the SB,
${\Psi}_{j,\delta}^{a}={\psi}_{j}^{\dagger}V^{a}_{\delta}{\psi}_{j+\delta},$ (61)
where $a=S,T,A,B$ refers to the four link operators introduced in Eq. (15), internal vertex function takes the following form:
$V^{\alpha}_{k,k+q}={1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{a^{\prime}a}_{\delta}\left((V_{\delta}^{a^{\prime}})^{\dagger}e^{-i{\bm{k}}\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{*}Pe^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}\right),$
where $\alpha$ denotes the component of fluctuation field $\Phi_{j,\delta}^{a}$ (see below for definition of matrix $P$), ${\bm{\delta}}$ is the bond vector of bond $\delta$,
$V^{\alpha}_{k,k+q}=-{1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{aa^{\prime}}_{\delta}\left(V_{\delta}^{a^{\prime}}e^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{T}Pe^{-i{\bm{k}}\cdot{\bm{\delta}}}\right),$
where $\alpha$ denotes the component of fluctuation field $\bar{\Phi}_{j,\delta}^{a}$, and
$V^{\alpha}_{k,k+q}={i\over 2}I_{2N}.$ (64)
where $\alpha\equiv\lambda$. Note that, to simplify the diagrammatic calculation, we have symmetrized the internal vertex function by implementing the particle-hole symmetry of the Nambu field, $\psi_{k}=P(\psi_{-k}^{\dagger})^{T}$, with $P=(\sigma_{x}\otimes I_{N})$.
Figure 3(a) and 3(b) illustrate the building blocks for a diagrammatic representation based on the formulation described above. We note that, since the Nambu fields $\Phi_{-q}$ and $\bar{\Phi}_{q}$ refer to the same physical degree of freedom, one can rewrite $S_{1}$ in an equivalent form:
$S_{1}\!=\!\!\frac{1}{2\sqrt{{\cal N}_{D}\beta}}\sum_{kq}\psi_{k}^{\dagger}\left(\delta_{q,0}G_{\rm sp}^{-1}(k)+2\sum_{\alpha}\bar{\Phi}^{\bar{\alpha}}_{q}V^{\alpha}_{k,k+q}\right)\psi_{k+q},$
where $\bar{\alpha}$ is defined such that $\bar{\Phi}^{\bar{\alpha}}_{-q}\equiv\Phi_{q}^{\alpha}$. Diagrammatically, this corresponds to an internal vertex with an in-coming wavy line (propagator of the fluctuation field, defined below), while that in Eq. (57) to an internal vertex with an out-coming wavy line.
In the last step, one can integrate out the boson fields, which yields
${\cal Z}=\int{\cal D}[\bar{\Phi}\Phi]\exp\left[-{N\over 4}S_{\rm eff}(\bar{\Phi},\Phi)\right],$
where
$S_{\rm eff}(\bar{\Phi},\Phi)={1\over 2}\sum_{q,\alpha,\alpha^{\prime}}\bar{\Phi}_{q}^{\alpha}{\Pi}_{0}^{\alpha\alpha^{\prime}}\Phi_{q}^{\alpha^{\prime}}+{2\over N}\text{Tr}[\ln{\cal M}]+{4S_{\rm cl}\over N}.$
In the limit of $N\to\infty$, the fluctuations of the auxiliary field $\Phi$ are suppressed, and only the SP action contributes to the partition function (namely, the SP approximation). For finite $N$, the typical amplitude of fluctuations of the field $\Phi$ is of order ${\cal O}(1/\sqrt{N})$, which enables us to expand the second term of $S_{\rm eff}$ as follows:
${2\over N}\text{Tr}[\ln{\cal M}]={2\over N}\text{Tr}\ln G_{\rm sp}^{-1}+{2\over\sqrt{N}}\sum_{m=1}^{\infty}{(-1)^{m+1}\over m}\text{Tr}\left[{\mathscr{V}}^{m}\right],$ (68)
where the trace “${\rm Tr}$” runs over $k$ and the Nambu indices, ${G^{-1}_{\rm sp}}$ is block diagonal in $k$, with each block given by $G^{-1}_{\rm sp}(k)$, while the matrix ${\mathscr{V}}$ is defined as
$\mathscr{V}_{k,k+q}=G_{\rm sp}(k){2\over\sqrt{{\cal N}_{D}\beta}}\sum_{\alpha}\Phi^{\alpha}_{-q}V^{\alpha}_{k,k+q}.$ (69)
In this expansion, Eq. (68), the first term is a constant and only contributes to static properties, like the ground state energy; the linear contribution in $\Phi$ vanishes for an expansion around a SP of the effective action; the quadratic contribution in $\Phi$ is combined with the first term of Eq. (67) to obtain the Gaussian contribution, $S^{(2)}(\bar{\Phi},\Phi)$, to the expansion of the effective action in powers of the fluctuations of the auxiliary fields. $S^{(2)}(\bar{\Phi},\Phi)$ determines the “bare” propagator of the fluctuation fields, which is usually known as “random phase approximation” (RPA) propagator. Thus, we reorganize the effective action as
$S_{\rm eff}(\bar{\Phi},\Phi)$ $=$ $\!{4S_{\rm cl}\over N}\!+\!{2\over N}\text{Tr}\ln G_{\rm sp}^{-1}\!+\!S^{(2)}(\bar{\Phi},\Phi)\!+\!S_{\rm int}(\bar{\Phi},\Phi),$
where
$S^{(2)}(\bar{\Phi},\Phi)={1\over 2}\sum_{q,\alpha,\beta}\bar{\Phi}_{q}^{\alpha}({\Pi}_{0}^{\alpha\alpha^{\prime}}-\Pi^{\alpha\alpha^{\prime}}(q))\Phi_{q}^{\alpha^{\prime}},$ (71)
is the Gaussian action of the auxiliary fields and
$\Pi^{\alpha\alpha^{\prime}}(q)={8\over N{\cal N}_{D}\beta}\sum_{k}{\rm tr}\left(G_{\rm sp}(k)V^{\bar{\alpha}}_{k,k+q}G_{\rm sp}(k+q)V^{\alpha^{\prime}}_{k+q,k}\right),$
is the polarization bubble illustrated in Fig. 3(d), where the trace “${\rm tr}$” runs over the Nambu indices. The polarization bubble determines the “bare” or RPA propagator of the auxiliary fields,
$D_{\alpha\alpha^{\prime}}(q)$ $=$ ${\int{\mathcal{D}}[\bar{\Phi},\Phi]e^{-{N\over 4}S^{(2)}(\bar{\Phi},\Phi)}\Phi_{\alpha}(q)\bar{\Phi}_{\alpha^{\prime}}(q)\over\int{\mathcal{D}}[\bar{\Phi},\Phi]e^{-{N\over 4}S^{(2)}(\bar{\Phi},\Phi)}}$ (73)
$=$ ${4\over N}\left[(\Pi_{0}-\Pi(q))^{-1}\right]_{\alpha\alpha^{\prime}}.$
In absence of a condensate, the RPA propagator scales as $1/N$ and is represented by the wavy line shown in Fig. 3(d). The last term of Eq. (LABEL:eq:Seffexp),
$S_{\rm int}(\bar{\Phi},\Phi)={2\over N}\sum_{m\geq 3}^{\infty}{(-1)^{m+1}\over m}\text{Tr}\left[{\mathscr{V}}^{m}\right],$ (74)
describes the effective interaction between the fluctuation fields. As shown in Fig. 4 for $m\leq 6$, the effective interaction vertex generated by each term of the expansion is diagrammatically represented as an “internal loop” of $m$ bosonic propagators and $m$ internal vertices.
Figure 4: Internal loops from the first four terms in $S_{\text{int}}$, Eq. (74).
In absence of a SB condensate, the power counting of each diagram, $1/N^{\nu}$, is determined by the difference between the number of RPA propagators ($P$) and the number of internal loops ($L$): $\nu=P-L$. As an example, Fig. 5 shows all the diagrams up to order $1/N$ of the magnetic susceptibility (see Sec. V for definition and Appendix A for details of the large-$N$ expansion of the magnetic susceptibility). The leading order diagram, $\nu=P-L=0$ because $P=L=0$, is shown in Fig. 5 (a). Fig. 5 (b) and (c) show the diagrams of order $1/N$. However, when a fraction of the bosons condense, a subset of these diagrams violates this nominal power counting. Specifically, diagrams that are nominally of order $1/N$ can acquire singular contributions of order $1/N^{0}$, which are essential for producing physically sound results [47, 48, 49].
Figure 5: Large-$N$ expansion of the magnetic susceptibility: (a) leading order diagram, equivalent to SP approximation, (b,c) sub-leading order diagrams in $1/N$. Notation of propagators and vertex functions are explained in Fig. 3.
In the following sections, we first describe the SP solution for both the QPM and AFM phases. We then analyze the $1/N$ corrections using the framework described in this section.
IV Saddle point solution
We consider the following ansatz for the SP solution of the link fields:
$\langle{S}_{j,\delta}\rangle\rvert_{\rm sp}$ $=$ ${\cal S},\quad\langle{T}_{j,\delta}\rangle\rvert_{\rm sp}={\cal T},$
$\langle{A}_{j,\delta}\rangle\rvert_{\rm sp}$ $=$ ${\cal A},\quad\langle{B}_{j,\delta}\rangle\rvert_{\rm sp}={\cal B},\quad\lambda_{j}\rvert_{\rm sp}=-i\tilde{\lambda}.$
which is motivated by symmetry considerations. More specifically, the QPM phase results from the condensation of only the singlet boson, with the corresponding gapless singlet excitation at momentum ${\bm{k}}={\bm{0}}$. Note that in our $SU(4)$ theory the singlet boson has dynamics, in contrast to the conventional bond-operator theory. For a proper gauge choice of the condensate wavefunction, both $\langle{S}_{j,\delta}\rangle\rvert_{\rm sp}$ and $\langle{T}_{j,\delta}\rangle\rvert_{\rm sp}$ are real numbers and transform trivially under the space group. The expectation values of link fields $\langle{A}_{j,\delta}\rangle\rvert_{\rm sp}$ and $\langle{B}_{j,\delta}\rangle\rvert_{\rm sp}$ are determined by the coupling with fields ${S}_{j,\delta}$ and ${T}_{j,\delta}$, i.e., by Eq. (26). Since the resulting SP Hamiltonian must retain the symmetries of the original model, $\langle{A}_{j,\delta}\rangle\rvert_{\rm sp}$ and $\langle{B}_{j,\delta}\rangle\rvert_{\rm sp}$ must transform like $\langle{S}_{j,\delta}\rangle\rvert_{\rm sp},\langle{T}_{j,\delta}\rangle\rvert_{\rm sp}$ under the symmetry group of $\hat{\mathscr{H}}$. The SP solution for the Lagrange multiplier, ${\tilde{\lambda}}_{j}\rvert_{\rm sp}={\tilde{\lambda}}$, is independent to site index $j$ because of translation invariance.
To determine the value of the above SP parameters, we need to solve the SP equations in Eq. (III.2). On the right-hand side of the SP equations, the average of the link operators and the boson density depend on the Green’s function of the SBs for a given SP configuration of the auxiliary fields, which is defined by the inverse of the dynamic matrix in Eq. (58). Essentially, this SP solution is equivalent to solving the eigenvalues and eigenstates of the following SP Hamiltonian in the canonical formulation of the problem
^$\hat{{\mathscr{H}}}_{\rm sp}={1\over 2}\sum_{\bm{k}}\hat{\psi}_{\bm{k}}^{\dagger}H_{\rm sp}({\bm{k}})\hat{\psi}_{\bm{k}},$ (76)
where $H_{\rm sp}({\bm{k}})$ has been defined in Eq. (58) and $\hat{\psi}_{\bm{k}}$ are the Nambu operators of the SBs in canonical formulation (c.f. Eq. (60)). Given the conservation of total spin, the SP Hamiltonian $H_{\rm sp}({\bm{k}})$ is block diagonal in the flavor of SBs, with each block given by
$H_{{\rm sp},\mu}({\bm{k}})$ $=$ $\left(\begin{array}[]{cc}A_{{\bm{k}},\mu}&B_{{\bm{k}},\mu}\\ B_{{\bm{k}},\mu}&A_{{\bm{k}},\mu}\end{array}\right)$ (79)
in the Nambu basis $(\hat{b}_{{\bm{k}},\mu},\hat{b}^{\dagger}_{-{\bm{k}},\mu})^{T}$. The matrix elements are:
$A_{{\bm{k}},0}$ $=$ $\tilde{\lambda}-J+J^{\prime}{\cal B}\gamma_{{\bm{k}}},\quad B_{{\bm{k}},0}=J^{\prime}{\cal A}\gamma_{{\bm{k}}},$
$A_{{\bm{k}},1}$ $=$ $A_{{\bm{k}},2}=A_{{\bm{k}}+{\bm{\pi}},3}=\tilde{\lambda}-J^{\prime}({\cal T}+{\cal B})\gamma_{{\bm{k}}},$
$B_{{\bm{k}},1}$ $=$ $B_{{\bm{k}},2}=B_{{\bm{k}}+{\bm{\pi}},3}=-J^{\prime}({\cal S}-{\cal A})\gamma_{{\bm{k}}},$ (80)
and
$\gamma_{\bm{k}}=\cos{k_{x}}+\cos{k_{y}}.$ ) .
The diagonal form of $H_{{\rm sp},\mu}({\bm{k}})$ is obtained by applying the following Bogoliubov transformation for each flavor:
$\left(\begin{array}[]{c}\hat{b}_{{\bm{k}},\mu}\\ \hat{b}_{\bar{\bm{k}},\mu}^{\dagger}\end{array}\right)=\left(\begin{array}[]{cc}u_{{\bm{k}},\mu}&v_{{\bm{k}},\mu}\\ v_{{\bm{k}},\mu}&u_{{\bm{k}},\mu}\end{array}\right)\left(\begin{array}[]{c}\hat{\beta}_{{\bm{k}},\mu}\\ \hat{\beta}_{\bar{\bm{k}},\mu}^{\dagger}\end{array}\right),$ (87)
where $\bar{\bm{k}}\equiv-{\bm{k}}$,
$u_{{\bm{k}},\mu}$ $=$ $\sqrt{{1\over 2}\left({A_{{\bm{k}},\mu}\over\varepsilon_{{\bm{k}},\mu}}+1\right)},$
$v_{{\bm{k}},\mu}$ $=$ $-\sqrt{{1\over 2}\left({A_{{\bm{k}},\mu}\over\varepsilon_{{\bm{k}},\mu}}-1\right)}\text{sgn}(B_{{\bm{k}},\mu})$ (88)
are coherent factors, and
$\varepsilon_{{\bm{k}},\mu}=\sqrt{A^{2}_{{\bm{k}},\mu}-B^{2}_{{\bm{k}},\mu}}$ (89)
is the SP energy spectrum.
Using the above solution, the Green’s function of the SBs can be computed straightforwardly by using Lehmann’s representation,
$G_{\rm sp}(k)$ $=$ $\frac{1}{H_{\rm sp}({\bm{k}})-i\nu_{m}}$ (90)
$=$ $\sum_{\varepsilon_{{\bm{k}},\mu}>0}\frac{g_{{\bm{k}},\mu}}{\varepsilon_{{\bm{k}},\mu}-i\nu_{m}}+\frac{{\bar{g}}_{{\bm{k}},\mu}}{\varepsilon_{{\bm{k}},\mu}+i\nu_{m}},$
where $g_{{\bm{k}},\mu}$ and ${\bar{g}}_{{\bm{k}},\mu}$ are $8\times 8$ matrices with non-zero matrix elements only for the entries corresponding to Nambu spinor $(\hat{b}_{{\bm{k}},\mu},\hat{b}^{\dagger}_{-{\bm{k}},\mu})^{T}$, given by
$g_{{\bm{k}},\mu}$ $=$ ($\ \hline\cr&u_{{\bm{k}},\mu}^{2}&v_{{\bm{k}},\mu}u_{{\bm{k}},\mu}&\\ &v_{{\bm{k}},\mu}u_{{\bm{k}},\mu}&v_{{\bm{k}},\mu}^{2}&\\ \hline\cr&&&\end{array}\right),$ (95)
${\bar{g}}_{{\bm{k}},\mu}$ $=$ ($\ \hline\cr&v_{{\bm{k}},\mu}^{2}&v_{{\bm{k}},\mu}u_{{\bm{k}},\mu}&\\ &v_{{\bm{k}},\mu}u_{{\bm{k}},\mu}&u_{{\bm{k}},\mu}^{2}&\\ \hline\cr&&&\end{array}\right).$ (100)
It is important to note that, the above formula can be safely used for calculations when the system size is finite or when there is no condensation of bosons. A more general representation is described at the end of Sec. IV.
We can now compute the expectation values of the link operators and boson density in the SP equation (III.2), which gives rise to the following self-consistent equations:
${\cal S}$ $=$ ${1\over{\cal N}_{D}}\sum_{{\bm{k}}}u_{{\bm{k}},0}v_{{\bm{k}},0}\cos\left({{\bm{k}}\cdot{\bm{\delta}}}\right),$
${\cal T}$ $=$ ${1\over{\cal N}_{D}}\sum_{{\bm{k}}}v_{{\bm{k}},0}^{2}\cos\left({{\bm{k}}\cdot{\bm{\delta}}}\right),$
${\cal A}$ $=$ ${1\over{\cal N}_{D}}\sum_{{\bm{k}}}\sum_{\mu=1}^{3}\eta_{\mu}u_{{\bm{k}},\mu}v_{{\bm{k}},\mu}\cos\left({{\bm{k}}\cdot{\bm{\delta}}}\right),$
${\cal B}$ $=$ ${1\over{\cal N}_{D}}\sum_{{\bm{k}}}\sum_{\mu=1}^{3}\eta_{\mu}v^{2}_{{\bm{k}},\mu}\cos\left({{\bm{k}}\cdot{\bm{\delta}}}\right),$
11 11 $=$ ${1\over{\cal N}_{D}}\sum_{{\bm{k}}}\sum_{\mu=0}^{3}v^{2}_{{\bm{k}},\mu},$ (101)
where $\eta_{\mu}=(-1,-1,1)$ for $\mu=1,2,3$ and ${\bm{\delta}}=(1,0)$ or $(0,1)$ refers to the two nearest neighbor bonds of the square lattice of dimers. Since the system preserves the four-fold rotation symmetry, the above equations do not depend on the choice of ${\bm{\delta}}$.
Figure 6: (a) SP spectrum of singlet and triplet bosons in the lab reference frame, where the triplet boson spectrum is three-fold degenerate. The color scale follows panel (b), indicating different values of $g$ between $g=0$ (brown) and $g=g_{c}\simeq 0.42$ (blue). In the twisted reference frame, the triplet boson dispersions for bosons $\hat{b}_{j,1}$ and $\hat{b}_{j,3}$ are shifted by ${\bm{\pi}}$, so they are gapless at the $\Gamma$ point. (b) Minimal energy gap of triplet boson as a function of $g$. (c) Total condensate fraction $n_{c}$. (d) Fraction of singlet bosons in the condensate, $n_{c,{\bm{0}}}/n_{c}$, and SP magnetization, $\rvert\langle S_{j,\pm}^{x}\rangle_{\rm sp}\rvert=\sqrt{n_{c,\bf 0}n_{c\bf\pi}}$, as a function of $g$. Vertical black line in (b-d) marks the value of $g_{c}$ obtained from the SP solution.
Fig. 6 (a) shows the single boson dispersion for the singlet and triplet channels. The dispersions of the three triplet bosons ($\mu=1,2,3$), presented in the laboratory reference frame, are identical because the global $\mathrm{SU}(2)$ symmetry of the Hamiltonian is preserved by the QPM phase. The singlet boson dispersion is gapless at the $\Gamma$ point, where the occupation number becomes macroscopic at zero temperature. The triplet-boson dispersion is gapped for small $g$ values and it has a minimum at ${\bm{k}}={\bm{\pi}}$. As illustrated in Fig. 6 (a) and 6 (b), this triplet mode softens as $|g-g_{c}|^{\nu}$ when $g$ approaches a critical point $g_{c}\approx 0.42$. This softening signals a phase transition to a state exhibiting magnetic long-range order. Notably, this critical value aligns closely with the value of $g_{c}\approx 0.3984$ obtained from QMC simulations [14]. It is worth mentioning that a similar approach based on the more conventional $\mathrm{SU}(2)$ SB theory results in a critical value $g_{c}\approx 0.235$ [40, 50, 38], which significantly deviates from the QMC result. Interestingly, this critical value closely aligns with the value $g_{c}=0.25$ that is obtained from a semi-classical treatment based on $\mathrm{SU}(4)$ coherent states. It should be stressed that the gap closes as $|g-g_{c}|^{\nu}$, with a critical exponent $\nu=1$. This value, characteristic of a large-$N$ theory  [13], differs from $\nu\simeq 0.71$ predicted by QMC [15]. A more precise estimation of $\nu$ requires the inclusion of higher order corrections in $1/N$.
Before delving into the SP solution for the broken-symmetry state, it is worthwhile to compare our SP solution for the QPM with the ’bond operator’ approach introduced by Sachdev and Bhatt [30]. The approximation proposed by Sachdev and Bhatt can be derived from our large-$N$ expansion if we approximate the expectation values of the ${S}_{j,\delta}$ and ${T}_{j,\delta}$ link fields in Eq. (75) in the following way:
$\langle_{j,\delta}\rangle_{\rm sp}\approx\langle b_{j,0}\rangle_{\rm sp}\langle b_{j+\delta,0}\rangle_{\rm sp},$ $\langle_{j,\delta}^{\;}\rangle_{\rm sp}\approx\langle b^{\dagger}_{j+\delta,0}\rangle_{\rm sp}\langle b_{j,0}\rangle_{\rm sp},$
where we have written the equations in the canonical formalism to make direct contact with Ref. [30]. This approximation, which is characteristic of semi-classical treatments ($1/M$ expansion), removes the expansion parameter from the formalism (it is neither a $1/M$ nor a $1/N$ expansion). This is the basic reason why their approach does not capture correctly the Goldstone modes of the broken symmetry state. On the other hand, it leads to a triplon gap that closes as $|g-g_{c}|^{\nu}$ with $\nu=1$, characteristic of large-$N$ fixed point [13, 34], in contrast to the large-$M$ (generalized spin wave theory) prediction $\nu=1/2$. This observation reveals the “hidden” large-$N$ nature of the original BOT mean-field approximation.
IV.1 Broken symmetry phase
The SP dispersion of the triplon modes remains gapless beyond the critical point. By evaluating the occupation number of the triplet bosons at the gapless momenta (${\bm{\pi}}$ in the laboratory frame), we find the density of triplet bosons that are condensed. We label the condensate fraction of singlet and triplet bosons as $n_{c,{\bm{0}}}$ and $n_{c,{\bm{\pi}}}$, respectively, and the total condensate fraction as $n_{c}=n_{c,{\bm{0}}}+n_{c,{\bm{\pi}}}$. As shown in Fig. 6 (c), the total condensate density ($n_{c}$) decreases with increasing $g$. The fraction $n_{c,{\bm{0}}}/n_{c}$ of singlet bosons in the condensate is shown in Fig. 6 (d).
When the triplet boson undergoes condensation, the system is generally expected to exhibit long-range Néel magnetic order. However, within the SP solution framework discussed above, the ground state expectation value of local spin operators remains zero because the SP solution retains the $\mathrm{SU}(2)$ invariance of the Hamiltonian. To accurately describe the condensate corresponding to a magnetically ordered state, it is necessary to introduce an infinitesimal symmetry-breaking field (SBF) [45], which is sent to zero after taking the thermodynamic limit.
In our current scenario, the SBF is a staggered magnetic field $h$ linearly coupled to the local order parameter. In the twisted reference frame, the corresponding Zeeman term takes the form:
^$\hat{{\mathscr{H}}}_{\rm SBF}=-h\sum_{j}(S_{j,+}^{x}-S_{j,-}^{x})=-h\sum_{j}(\hat{b}_{j,0}^{\dagger}\hat{b}_{j,1}+\hat{b}_{j,1}^{\dagger}\hat{b}_{j,0}).$ (103)
In the thermodynamic limit, the magnitude of this SBF is infinitesimally small, thereby leaving the SP spectrum unchanged. However, this field significantly alters the condensate wavefunction by controlling the relative population of bosons in the four gapless states. When solving the SP equation under the influence of this SBF, we find that all the bosons condense into a hybridized energy level formed by the singlet boson ($\mu=0$) and one flavor ($\mu=1$) of the triplet boson. The resulting condensate maintains the same fractions, $n_{c,{\bm{0}}}$ and $n_{c,{\bm{\pi}}}$, as those obtained without the SBF, which is essential for satisfying the SP equation because the contribution from non-condensed bosons remains unchanged. Consequently, this condensate induces Néel magnetic order polarized along the $x$ direction, with the magnitude of the ordered moment determined by $\sqrt{n_{c,{\bm{0}}}n_{c,{\bm{\pi}}}}$.
To understand how the presence of the SBF modifies the condensate, we consider the SP Hamiltonian for the condensed bosons under the SBF,
^$\hat{{\mathscr{H}}}_{\rm sp,c}=\frac{1}{2}\hat{\psi}_{01}^{\dagger}H_{01}\hat{\psi}_{01}+\frac{1}{2}\hat{\psi}_{2}^{\dagger}H_{2}\hat{\psi}_{2}+\frac{1}{2}\hat{\psi}_{3}^{\dagger}H_{3}\hat{\psi}_{3},$
where $\hat{\psi}_{01}^{\dagger}\equiv(\hat{b}_{\bm{0},0}^{\dagger},\hat{b}_{\bm{0},1}^{\dagger},\hat{b}_{\bm{0},0},\hat{b}_{\bm{0},1})$, $\hat{\psi}_{2}^{\dagger}\equiv(\hat{b}_{{\bm{0}},2}^{\dagger},\hat{b}_{{\bm{0}},2})$, $\hat{\psi}_{3}^{\dagger}\equiv(\hat{b}_{{\bm{\pi}},3}^{\dagger},\hat{b}_{{\bm{\pi}},3})$, and the Hamiltonian matrices are
$H_{01}$ $=$ $\left(\begin{array}[]{cccc}\Delta_{0}+\delta_{0}&-h&\Delta_{0}&0\\ -h&\Delta_{3}+\delta_{3}&0&\Delta_{3}\\ \Delta_{0}&0&\Delta_{0}+\delta_{0}&-h\\ 0&\Delta_{3}&-h&\Delta_{3}+\delta_{3}\end{array}\right),$ (109)
$H_{2}$ $=$ $H_{3}=\left(\begin{array}[]{cc}\Delta_{3}+\delta_{3}&\Delta_{3}\\ \Delta_{3}&\Delta_{3}+\delta_{3}\end{array}\right),$ (112)
with $\Delta_{0}=2J^{\prime}{\cal A}$ and $\Delta_{3}=-2J^{\prime}({\cal S}-{\cal A})$. The parameters $\delta_{0}$ and $\delta_{3}$ represent the corrections to the SP Hamiltonian due to the finite $h$. In the thermodynamic limit, one can verify that the four energies reduce to zero when $\delta_{0}=\delta_{3}=0$, corresponding to $h=0$. This implies that, in absence of the SBF $h$, the ground state is a non-magnetic condensate as the three triplet states exhibit identical condensate densities.
For finite $h$, the spectrum of the hybridized singlet- ($\mu=1$) triplet bosons, determined by $H_{01}$, becomes
$E_{0}=\sqrt{a-b},\quad E_{1}=\sqrt{a+b},$ (113)
where
$a$ $=$ $\Delta_{0}\delta_{0}+\Delta_{3}\delta_{3}+\frac{1}{2}(\delta_{0}^{2}+\delta_{3}^{2})+h^{2},$ (114)
$b$ $=$ $\frac{1}{2}\left[\left(2\Delta_{0}\delta_{0}+2\Delta_{3}\delta_{3}+\delta_{0}^{2}+\delta_{3}^{2}+h^{2}\right)^{2}\right.$
-- $\left.4\left((2\Delta_{0}+\delta_{0})(2\Delta_{3}+\delta_{3})-h^{2}\right)(\delta_{0}\delta_{3}-h^{2})\right]^{1/2}.$
The spectrum of $H_{2}$ and $H_{3}$ reads
$E_{2}=E_{3}=\sqrt{2\Delta_{3}\delta_{3}+\delta_{3}^{2}}.$ (116)
To determine the new ground state, we expand $\delta_{0}$ and $\delta_{3}$ in powers of $h$: $\delta_{0}=a_{0}h+{\cal O}(h^{2})$ and $\delta_{3}=a_{3}h+{\cal O}(h^{2})$. By inserting these expressions in Eqs. (LABEL:eq:hybH) and (116), we obtain the expressions for the squares of the four energies to linear order in $h$:
$E_{0}^{2}$ $\simeq$ $\left[\Delta_{0}a_{0}+\Delta_{3}a_{3}-\sqrt{(\Delta_{0}a_{0}-\Delta_{3}a_{3})^{2}+4\Delta_{0}\Delta_{3}}\right]h,$
$E_{1}^{2}$ $\simeq$ $\left[\Delta_{0}a_{0}+\Delta_{3}a_{3}+\sqrt{(\Delta_{0}a_{0}-\Delta_{3}a_{3})^{2}+4\Delta_{0}\Delta_{3}}\right]h,$
$E_{2}^{2}$ $=$ $E_{3}^{2}\simeq 2\Delta_{3}a_{3}h.$ (117)
The requirement of a SP solution with a finite condensate fraction necessitates that $E_{0}=0$, which in turn leads to the condition $a_{0}a_{3}=1$. This condition implies that the other three modes have excitation energies proportional to $\sqrt{h}$, so the bosons condense only in the single-particle state associated with the $E_{0}$ mode.
The ratio $a_{0}/a_{3}$ is the determined from the condition that the ground state for infinitesimal field $h$ must preserve the ratio $n_{c,{\bm{0}}}/n_{c,{\bm{\pi}}}$ between the condensate fractions $n_{c,{\bm{0}}}$ and $n_{c,{\bm{\pi}}}$ in the singlet and triplet states.
The wavefunction associated with the hybridized level $E_{0}$,
$X_{c}=\sqrt{{\cal N}_{D}}\left(\sqrt{n_{c,{\bm{0}}}},\sqrt{n_{c,{\bm{\pi}}}},0,0,-\sqrt{n_{c,{\bm{0}}}},-\sqrt{n_{c,{\bm{\pi}}}},0,0\right)^{T},$ (118)
is given in the Nambu representation for SB, as described in Eq. (60) for $N=4$ specifically. A normalization factor $\sqrt{{\cal N}_{D}}$ is introduced to account for finite size dimer lattices.
This solution explicitly breaks the $\mathrm{SU}(2)$ spin rotation symmetry and thus gives rise to local magnetic moments in the ground state, which are obtained by replacing the boson operator $\hat{b}_{j,\mu}$ with its SP expectation value: $\langle b_{j,\mu}\rangle_{\rm sp}=(\sqrt{n_{c,{\bm{0}}}},\sqrt{n_{c,{\bm{\pi}}}},0,0)$. As expected, this condensate corresponds to an AFM state polarized along the $x$ direction, with magnitude of the ordered moment:
$\langle S^{x}_{j,+}\rangle_{\rm sp}=-\langle S^{x}_{j,-}\rangle_{\rm sp}=\sqrt{n_{c,{\bm{0}}}n_{c,{\bm{\pi}}}}.$ (119)
Figure 7: SP solution of the antisymmetric DSSF $S^{\mu\mu}_{{\rm sp},A}({\bm{q}},\omega)$ for different values of $g$, where $S^{xx}_{{\rm sp},A}({\bm{q}},\omega)=S^{yy}_{{\rm sp},A}({\bm{q}},\omega)=S^{zz}_{{\rm sp},A}({\bm{q}},\omega)$ is isotropic in QPM phase, and $S^{yy}_{{\rm sp},A}({\bm{q}},\omega)=S^{zz}_{{\rm sp},A}({\bm{q}},\omega)$ refers to the transverse spin components in AFM phase. For visual clarity, we represent the $\delta$-peak in the SB result with thick dark red lines. The white circles indicate the energy dispersion obtained from QMC simulations.
For a finite condensate fraction, the SP Green’s function becomes singular for momenta with a gapless spectrum. This singularity can be regularized by considering a finite-size system, as is done when discussing the SP Green’s function in Eq. (90) and the SP equations in Eq. (IV). Alternatively, one can isolate the singular part and then take the thermodynamic limit, which separates the SP Green’s function into two components:
$G_{\rm sp}(k)=G_{c}(k)+G_{n}(k).$ (120)
The first term on the right-hand side is the SP Green’s function of the condensed SBs,
$G_{c}(k)=g_{c}(2\pi)^{3}\delta({\bm{k}})\delta(\omega_{m}),$ (121)
with
$g_{c}=\lim_{{\cal N}_{D}\to\infty}{1\over{\cal N}_{D}}X_{c}X_{c}^{\dagger},$ (122)
and $X_{c}$ given by Eq. (118). The second term is the SP Green’s function of the non-condensed SBs, which takes the same form as that in Eq. (90). We note that, the above expression for the SP Green’s function applies to both the QPM phase, for which $n_{c,{\bm{\pi}}}=0$, and the AFM phase.
V Magnetic excitation spectrum
Figure 8: SP result of the static structure factor $S^{zz}_{{\rm sp},A}({\bm{q}})$ defined in Eq. (134) for different values of $g$. The structure factor is independent of the $\mu$ component in the isotropic QPM phase. For the AFM phase, $S^{zz}_{{\rm sp},A}({\bm{q}})=S^{yy}_{{\rm sp},A}({\bm{q}})$ correspond to the two transverse spin components. The blue lines refer to the SP result, while black circles are obtained from the QMC simulations.
Having solved the SP equations, we are now equipped to study the magnetic excitations described by the imaginary part of the magnetic susceptibility (fluctuation-dissipation theorem). Within the path-integral framework, the magnetic susceptibility is defined by the second derivative of the free energy with respect to an external time-space dependent magnetic field:
$\chi^{\mu\nu}_{\sigma\sigma^{\prime}}({\bm{q}},i\omega_{m})=\frac{\delta^{2}\ln{\cal Z}(h)}{\delta h_{-q\sigma}^{\mu}\delta h_{q\sigma^{\prime}}^{\nu}},$ (123)
where $\mu,\nu=x,y,z$ denotes the spin component indices, $\sigma,\sigma^{\prime}=\pm$ label the two sites within a dimer. ${\cal Z}(h)$ is the partition function in the presence of the external magnetic field $h_{j,\pm}^{\mu}$ coupled to the local moments via the action:
[…]
Given the invariance of the system under a mirror operation that exchanges the two layers (up to a translation by one lattice space for the AFM phase), it is convenient to introduce the symmetric (S) and antisymmetric (A) combinations of spin operators within a dimer:
$u_{S}^{\mu}=u_{+}^{\mu}+u_{-}^{\mu},\ \ u_{A}^{\mu}=u_{+}^{\mu}-u_{-}^{\mu}.$ (125)
along with the simplified notation for the magnetic susceptibility, $\chi_{\alpha}^{\mu\nu}$, where the subscript $\alpha=S,A$ denotes the symmetric or antisymmetric channels, respectively. Due to the $\mathrm{SU}(2)$ rotation symmetry, or the residual $\mathrm{U}(1)$ symmetry about the local moments in the Néel order phase, the magnetic susceptibility is diagonal in the spin index. The diagonal components satisfy $\chi_{\alpha}^{xx}(q)=\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the QPM phase and $\chi_{\alpha}^{xx}(q)\neq\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the Néel ordered phase, assuming that the local moments align along the $x$ direction.
In the remainder of this work, we focus on the antisymmetric channel, as it carries the dominant spectral weight within the parameter range of interest. For completeness, a parallel discussion of the symmetric channel is included in Appendix C.
V.1 Saddle-point approximation
In the SP approximation, the magnetic susceptibility is given by the diagram shown in Fig. 5 (a):
$\chi_{{\rm sp},\alpha}^{\mu\mu}(q)=\frac{1}{2{{\cal N}_{D}}\beta}\sum_{k}{\rm tr}\left[G(k)u^{\mu}_{\alpha}G(k+q)u^{\mu}_{\alpha}\right].$ (126)
The susceptibility along the real-frequency axis is obtained through analytic continuation ($i\omega_{n}\to\omega+i0^{+}$). According to the fluctuation-dissipation theorem, the dynamic spin structure factor is proportional to the imaginary part of the susceptibility:
$S_{{\rm sp},\alpha}^{\mu\mu}({\bm{q}},\omega)=-\frac{1}{\pi}\text{Im}[\chi_{{\rm sp},\alpha}^{\mu\mu}({\bm{q}},\omega)].$ (127)
As shown in Fig. 7, the DSSF in the antisymmetric channel ($\alpha=A$) consists of a $\delta$-peak and a continuum of excitations. To understand the origin of these two contributions we can split $\chi_{{\rm sp},A}^{\mu\mu}$ into two parts, $\chi_{{\rm sp},A}^{\mu\mu}=\chi_{{\rm sp},A}^{\mu\mu(1)}+\chi_{{\rm sp},A}^{\mu\mu(2)}$, where
$\chi_{{\rm sp},A}^{\mu\mu(1)}({\bm{q}},\omega)$ $=$ $\frac{1}{2}{\rm tr}\left[g_{c}u^{\mu}_{A}G_{n}({\bm{q}},\omega+i0^{+})u^{\mu}_{A}\right]$ (128)
$+$ $\frac{1}{2}{\rm tr}\left[G_{n}(-{\bm{q}},-\omega-i0^{+})u^{\mu}_{A}g_{c}u^{\mu}_{A}\right]$
involves one condensed SB. This contribution has a single-boson pole arising from SP Green’s function. The contribution from the second term, $\chi_{{\rm sp},A}^{\mu\mu(2)}$, becomes apparent after performing the Matsubara frequency summation,
$\chi_{{\rm sp},A}^{\mu\mu(2)}({\bm{q}},\omega)=\frac{1}{2{{\cal N}_{D}}}\sum_{n,m=0}^{3}{\sum_{\bm{k}}}^{\prime}\frac{{\rm tr}\left[g_{{\bm{k}},n}u^{\mu}_{A}{\bar{g}}_{{\bm{k}}+{\bm{q}},m}u^{\mu}_{A}\right]}{\varepsilon_{{\bm{k}},n}+\varepsilon_{{\bm{k}}+{\bm{q}},m}+\omega+i0^{+}}+\frac{{\rm tr}\left[{\bar{g}}_{{\bm{k}},n}u^{\mu}_{A}g_{{\bm{k}}+{\bm{q}},m}u^{\mu}_{A}\right]}{\varepsilon_{{\bm{k}},n}+\varepsilon_{{\bm{k}}+{\bm{q}},m}-\omega-i0^{+}},$ (129)
where the prime symbol indicates that the sum is restricted to momenta ${\bm{k}}$ such that $\varepsilon_{{\bm{k}},n}>0,\varepsilon_{{\bm{k}}+{\bm{q}},m}>0$. This contribution gives rise to the two-particle continuum in the DSSF  [^3] In the SP approximation, the two-particle continuum corresponds to the excitation of a singlet boson and a triplet boson (see Eq. (130)). However, this is a spurious continuum, as the singlet boson is not a physical mode. The physical continuum appears at higher energies and corresponds to a three-triplon excitations, which are only obtained after properly accounting for quantum fluctuations (this a generic feature of the large-$N$ theories). In contrast, the two-particle continuum in the symmetric channel of the DSSF is physical, as it consists of two triplon excitations..
In the QPM phase, the $\delta$-peak in the DSSF coincides with the SP spectrum of the three triplet bosons (triplon modes). Physically, these modes are excited by the antisymmetric spin operator:
$\hat{S}_{jA}^{\mu}=\hat{S}_{j+}^{\mu}-\hat{S}_{j-}^{\mu}=\hat{b}_{j,0}^{\dagger}\hat{b}_{j,\mu}+\hat{b}_{j,\mu}^{\dagger}\hat{b}_{j,0},$ (130)
because singlets and triplets have opposite parity under the mirror symmetry that exchanges the two layers. In the presence of a singlet-boson condensate, $\langle\hat{b}_{j,0}\rangle\simeq\sqrt{n_{c}}$ (here we fix the overall complex phase of the condensate to make this value real), we can approximate
$\hat{S}_{jA}^{\mu}\simeq\sqrt{n_{c}}\left(\hat{b}_{j,\mu}+\hat{b}_{j,\mu}^{\dagger}\right),$ (131)
which, upon acting on the ground state, creates a specific flavor of the triplet boson, corresponding to the $\delta$-peak in the DSSF.
Figure 9: Longitudinal DSSF in the anti-symmetric channel for $g=0.8$. (a) SP result including two $\delta$-peaks corresponding to two unphysical modes. (b) Result after including the counter-diagram shown in Fig. 11 (b). The lines indicate the single-boson dispersion obtained from the SP Hamiltonian. The dotted square denotes the region where an extra low-intensity quasi-flat mode emerges from the fluctuations of the auxiliary fields. (c) Cancellation of the $\delta$-peak of the SP susceptibility upon adding the the counter-diagram shown in Fig. 11 (b). (d) Low-frequency scaling of the intensity of the continuum at the ordering wavevector $(\pi,\pi)$. The black dashed line indicates the slope associated with $1/\omega$ scaling.
In the AFM phase, both the singlet and the triplet boson parallel to the ordered moment condense, $\langle\hat{b}_{j,0}\rangle\sim\sqrt{n_{c,{\bm{0}}}}$ and $\langle\hat{b}_{j,1}\rangle\sim\sqrt{n_{c,\pi}}$ (up to a complex phase), while the other two transverse triplet bosons remain uncondensed. Consequently, the two transverse spin operators can be approximated by :
$\hat{S}_{jA}^{\mu}\simeq\sqrt{n_{c,{\bm{0}}}}\left(\hat{b}_{j,\mu}+\hat{b}_{j,\mu}^{\dagger}\right)\quad\text{for}\ \mu=y,z.$ (132)
This operator generates the corresponding flavor of triplet bosons upon acting on the ground state, corresponding to the magnon excitations of the system. In contrast, the longitudinal spin operator,
$\hat{S}_{jA}^{x}\simeq\sqrt{n_{c,{\bm{0}}}}\left(\hat{b}_{j,1}+\hat{b}_{j,1}^{\dagger}\right)+\sqrt{n_{c,{\bm{\pi}}}}\left(\hat{b}_{j,0}+\hat{b}_{j,0}^{\dagger}\right),$ (133)
generates two types of $\delta$-peaks arising from the $\hat{b}_{j,0}$ and $\hat{b}_{j,1}$ bosons, as shown in Fig. 9 (a). In the next section we will demonstrate that the $\delta$-peak arising from the $\hat{b}_{j,0}$ boson disappears upon adding the “counter-diagram” shown in Fig. 5 (b). Furthermore, the $\delta$-peak from the $\hat{b}_{j,1}$ boson gets broadened.
Remarkably, the DSSF described by the SP approximation of the $\mathrm{SU}(4)$ SB approach shows very good agreement with QMC simulations. As demonstrated in Fig. 7, both the triplon dispersion in the QPM phase and the magnon dispersion in the Néel order phase match the QMC results with high precision. This agreement persists deep inside the magnetically ordered phase. Moreover, the spectral weight carried by these excitations closely aligns with the QMC results. In Fig. 8, we compare the total spectral weight at a given ${\bm{q}}$, obtained from the equal-time correlation function in the antisymmetric channel
$S_{A}^{\mu\mu}({\bm{q}})=\int_{0}^{\infty}\frac{d\omega}{2\pi}S_{A}^{\mu\mu}({\bm{q}},\omega).$ (134)
We find that the ${\bm{q}}$ dependence of $S_{A}^{\mu\mu}({\bm{q}})$ closely follows the QMC results, though the value obtained from the SB approach is systematically slightly higher than the QMC value. This overestimation of the spectral weight is a known feature of the mean field SB theory and it arises from violations of the local constraint. For instance, in the $\mathrm{SU}(2)$ SB approach, the integrated spectral weight is overestimated by a factor of $3/2$ [45]. As we will see in the next section, the inclusion a $1/N$ contribution reduces the integrated spectral weight and further improves the agreement between the SB approach and QMC.
Figure 10: Each diagram shown in panel (a) must be accompanied by the counter diagram shown in panel (b).
Figure 11: (a) SP diagram and (b) counter-diagram for the magnetic susceptibility. The counter-diagram vanishes for all components of the magnetic susceptibility in the QPM phase, and for the transverse components in the AFM phase. Arrows in the plot indicate the particle flow.
V.2 Amplitude fluctuation in Néel AFM phase
In this section, we examine the longitudinal spin fluctuations of the AFM phase. Although the longitudinal mode is expected to be gapped, it acquires an intrinsic broadening due to its decay into pairs of transverse Goldstone modes (i.e., magnons). In a two-dimensional system, the infrared divergence associated with this decay process raises the question of whether this mode can be distinguished from the two-magnon continuum. This issue has been investigated through various approaches in the literature, including field-theoretic analyses of the effective low-energy ${\mathrm{O}}(3)$ model describing long-wavelength fluctuations of local magnetic moments [52, 53, 54], as well as QMC studies [16, 55]. These studies conclude that the amplitude mode is barely visible in the dynamic spin structure factor for the 2D ${\mathrm{O}}(3)$ model. Here, we revisit this problem using the $\mathrm{SU}(4)$ SB approach and demonstrate that the same conclusion holds.
We first recall that, in the SP approximation, the longitudinal spin fluctuation $S_{{\rm sp},A}^{xx}({\bm{q}},\omega)$ consists of two $\delta$-peaks: one arising from the longitudinal triplet boson $\hat{b}_{{\bm{q}},1}$ and the other one arising from the singlet boson $\hat{b}_{{\bm{q}},0}$ [see Fig. 9 (a)]. This qualitatively incorrect result is characteristic of SP solutions with condensed bosons. As it has been explained at length for $\mathrm{SU}(2)$ SBs [47, 48, 49], in presence of a condensate, each diagram must be accompanied by a counter diagram where one external vertex is replaced by a “tail” consisting of the RPA propagator attached to a bubble with an internal and an external vertex (see Fig. 10). Since this bubble represents a cross susceptibility between the spin-components, which transform like a vector under $\mathrm{SU}(2)$ rotations and the auxiliary fields, which are scalars, it vanishes in the QPM phase along with the counter-diagram of the SP diagram, shown in Fig. 11 (b). This is the reason why the SP spectrum is qualitatively correct on the QPM side of the phase diagram.
For the AFM phase, the residual symmetry group is $\mathrm{U}(1)$. Since the transverse spin-components still transform like a vector under this group, the counter-diagram still vanishes for the transverse components of the spin susceptibility. This is not true, however, for the longitudinal spin component that is a scalar of the residual symmetry group. The counter-diagram, whose nominal order is $1/N$, includes singular contributions of order $1/N^{0}$ that exactly cancel the residues of the two poles of $\chi_{{\rm sp},A}^{xx}({\bm{q}},\omega)$, eliminating the aforementioned pair of $\delta$-peaks.
Figure 12: Longitudinal antisymmetric DSSF at the ordering wave vector ${\bm{\pi}}$. (a) Extra peak induced by the inter-particle interaction. (b) Softening of the extra peak as reducing inter-dimer interaction.
As derived in Appendix A, the diagram in Fig. 11 (b) is given by
$\chi_{{\rm fl},A}^{\mu\nu}(q)=\sum_{\alpha,\alpha^{\prime}}\Lambda^{\mu,\alpha}_{A}(q)D_{\alpha\alpha^{\prime}}(q)\Lambda_{A}^{\mu\bar{\alpha^{\prime}}}(-q),$ (135)
where
$\Lambda_{A}^{\mu,\alpha}({q})=\frac{1}{{\cal N}_{D}\beta}\sum_{k}{\rm tr}\left[G(k)u_{A}^{\mu}G({k+q})v^{\alpha}_{k+q,k}\right]$
is the cross susceptibility between the spin-components and the auxiliary fields, and $D_{\alpha\alpha^{\prime}}(q)$ is the propagator of the fluctuation field, defined in Eq. (73). Fig. 9 (b) presents the longitudinal antisymmetric DSSF in the AFM phase after including this contribution for $g=0.8$. We observe that the two $\delta$-peaks in the SP result, shown in Fig. 11 (a), are removed, as explicitly shown in Fig. 9 (c) for an arbitrary momentum ${\bm{q}}$.
Figure 13: Equal-time longitudinal spin structure factor $S^{xx}_{A}(\bm{q})$ for $g=0.8$ in the SP approximation of SB approach and that including the proper $1/N$ corrections, and the QMC result.
With both $\delta$-peaks removed, the remaining spectral weight forms a structured continuum. Focusing on the low-frequency regime around the ordering wave vector, the intensity of the continuum scales as $1/\omega$ (see Fig. 9 (d)). This scaling is already evident at the SP level, as detailed in Appendix B. It results from the product of the density of states, which scales as $\omega$ (linear dispersion) and the square of the matrix element that scales like $1/\omega^{2}$. This behavior is consistent with previous studies using different analytical and numerical methods [52, 53, 54, 16, 55].
Moving away from the ordering wave vector, the longitudinal DSSF displays a broadened peak at energy $v_{L}\rvert{\bm{q}}\rvert$ with an intrinsic broadening $\Gamma_{\bm{q}}$, where ${\bm{q}}$ is the momentum measured from the ordering wave vector ${\bm{\pi}}$. The velocity $v_{L}$ approximately equals that of the triplet bosons, as seen in Fig. 9 (b). The shape of this peak fits well with the formula $\text{Im}[Z_{\bm{q}}/\sqrt{(v_{L}\rvert{\bm{q}}\rvert)^{2}-(\omega+i\Gamma_{\bm{q}})^{2}}]$, corresponding to a pole in the magnetic susceptibility in the lower half of the complex frequency plane, $z=v_{L}\rvert{\bm{q}}\rvert-i\Gamma_{\bm{q}}$. As ${\bm{q}}\to\bm{0}$, both the real and imaginary parts of this pole reduce to zero, while $Z_{\bm{q}}$ remains finite, thereby recovering the $1/\omega$ form of the DSSF.
A closer examination of panels (a) and (b) of Fig. 9 reveals a low intensity quasi-flat mode emerging from the continuum around energy $3J$. The feature, which is absent in the SP contribution is more clearly displayed by the constant ${\bm{q}}={\bm{\pi}}$ plot, shown in Fig. 12 (a). The SP result shows a regular continuum over the energy range of interest, while the inclusion of $\chi_{{\rm fl},A}^{xx}(q)$ introduces a peak, signaling the emergence of a resonance state due to fluctuations of the auxiliary fields. As the inter-dimer interaction decreases, this peak gradually shifts to lower energies and becomes less prominent near the critical point, where it is overshadowed by the high-intensity ($1/\omega$ scaling) of the low-frequency continuum (see Fig. 12 (b)). Notably, the energy scale of this mode is comparable to the amplitude fluctuations observed in QMC simulations at $\omega\approx 1.5J$. Additionally, the DSSF displays a broad peak around $\omega\sim 4J$, which is already present in the SP solution and undergoes only minor modifications due to quantum fluctuation effects. This broad structure also appears in QMC simulations at a similar energy scale [16].
Finally, in Fig. 13 we compare the equal-time longitudinal structure factor, defined in Eq. (134), with the QMC result for $g=0.8$. The significant improvement over the SP result stems from a more accurate treatment of the local constraint in Eq. (29), leading to a better fulfillment of the sum rule.
VI Concluding remarks
As was observed by Perelomov in the seventies [56], coherent states of Lie algebras provide the natural link between quantum and classical mechanics. In fact, $N$-level quantum mechanical systems admit multiple classical limits corresponding to different choices of $N$-dimensional representations of distinct Lie algebras. Naturally, for each specific choice of Hamiltonian, there is one classical limit that better approximates dynamics of the quantum system.
Spin systems naturally provide examples of $N$-level units, where the unit may consist of a single spin $(N=2S+1)$ or a set of spins, such as the dimer units considered in this work. Recent studies have highlighted multiple instances of realistic spin Hamiltonians where the classical limit based on coherent states of the completely symmetric irreps of $\mathrm{SU}(N)$, labelled by the integer index $M$, offers a more accurate approximation of the exact dynamics than the classical limit based on $\mathrm{SU}(2)$ coherent states [36, 43, 44, 35].
Since the classical theory becomes exact in the limit $M\to\infty$, quantum corrections for describing the spin dynamics for a finite $M$ irrep can be implemented via a $1/M$ expansion. This expansion is achieved by introducing a faithful representation of spin operators in terms of SU$(N)$ Schwinger bosons and accounting for the local constraint (29) via a generalization of the Holstein-Primakoff (HP) transformation. The $1/M$ expansion arises from the Taylor series expansion of the square root function inherent to the HP transformation. The resulting quadratic Hamiltonian in HP bosons corresponds to the generalized spin wave Hamiltonian, with the effects of non-quadratic terms systematically included order by order in $1/M$. The diagrammatic representation of this $1/M$ expansion leads to the so-called “loop expansion”, where the order of each diagram in powers of $1/M$ is equal to the number of loops [57, 58].
In contrast to the semi-classical large-$M$ expansions described above, one can also use the SB representation of spin operators as a basis for implementing a $1/N$ expansion, where the spin model is generalized to more general groups labeled by $N$. Just as standard semi-classical large-$S$ approaches rely on $\mathrm{SU}(2)$ coherent states, traditional large-$N$ approaches utilize $\mathrm{SU}(2)$ SBs to represent spin operators. However, since other Lie algebras may be more suitable for semi-classical expansions, it is natural to conjecture that these alternative algebras could also provide a better foundation for implementing a $1/N$ expansion. The results presented in this manuscript confirm this conjecture. The two different strategies, large-$M$ and large-$N$, are schematically shown in Figure 2.
To test the above-mentioned conjecture, in this work we developed a large-$N$ $\mathrm{SU}(4)$ Schwinger boson theory for the spin-$1/2$ bilayer square lattice Heisenberg antiferromagnet. The spin operators of each dimer are represented by four SBs, that create the coherent states that represent any quantum mechanical state of a given dimer. A key distinction from standard large-$N$ approaches based on $\mathrm{SU}(2)$ SBs [45] is that the $\mathrm{SU}(4)$ SBs fully capture the intra-dimer entanglement, while the link fields and their fluctuations generate inter-dimer entanglement. To perform a large-$N$ expansion, the original dimer problem with two antiferromagnetically coupled $\mathrm{SU}(2)$ spins on each site of the dimer has been generalized to two antiferromagnetically coupled SU(n) spins, with $n\geq 2$. Then, $\mathrm{SU}(N=n^{2})$ SBs have been used to include all quantum mechanical states of the generalized dimer in the manifold of $\mathrm{SU}(N)$ coherent states. Both the static and dynamical properties are significantly better described by the large-$N$ expansion based on $\mathrm{SU}(4)$ SBs than by the large-$M$ $\mathrm{SU}$(4) limit.
Furthermore, the theory introduces an expansion parameter ($1/N$), that cures the shortcomings of previous mean-field approximations based on bond operators [30, 59, 20, 25]. While alternative approaches based on expansions in the inverse system dimension [31, 32] address the same problem, they still produce values of $g_{c}$, which deviate significantly from the numerical (QMC) values. Similar limitations are also observed in large-$M$ expansions based on $\mathrm{SU}(4)$ coherent states or large-$N$ expansions based on $\mathrm{SU}(2)$ SBs [40, 50].
Achieving quantitative agreement with the exact value of $g_{c}$ is crucial for precisely describing real materials near the quantum critical point. Typically, the exact value of $g_{c}$ cannot be determined from QMC simulations due to the well-known sign problem that affects most frustrated Hamiltonians of interest. Without precise knowledge of $g_{c}$, it is challenging to overcome the quantitative limitations of the aforementioned methods by simply rescaling $g$.
Therefore, the remarkable accuracy of the large-$N$ method introduced in this work marks a significant advancement in modeling the static and dynamical properties of coupled antiferromagnetic dimers near continuous quantum phase transitions between magnetically ordered and paramagnetic states.
Acknowledgements.
We thank Martin Mourigal for useful discussions. This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences, and Engineering Division under award DE-SC-0018660. E.A.G. is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. Y. K. is supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Numbers JP20H00122 and JP22K03509. A part of numerical calculations were performed using the facilities of the Supercomputer Center, The Institute for Solid State Physics, The University of Tokyo.
Appendix A Large-$N$ expansion of DSSF
To compute the dynamic magnetic susceptibility within the path-integral framework, we introduce an external source term coupled to the local magnetic moment, as described by Eq. (124) in the main text. The Hubbard-Stratonovich transformation is then applied to the quartic interaction term, and the action is expanded around the saddle point (SP), similar to the procedure without the source term. It is important to note that the SP remains the same as in the absence of the source term. After integrating out the boson fields, the effective action in the presence of the source term is given by:
$S_{\rm eff}(\bar{\Phi},\Phi,h)$ $=$ $\frac{1}{2}\sum_{q,\alpha,\alpha^{\prime}}(\bar{\Phi}^{\dagger})_{q}^{\alpha}{\Pi}_{0}^{\alpha\alpha^{\prime}}\Phi_{q}^{\alpha^{\prime}}+\frac{2}{N}\text{Tr}[\ln{\cal M}(h)]$ (137)
$+\frac{4S_{\rm cl}}{N}.$
The external source term appears in the matrix ${\cal M}(h)$, whose matrix elements are defined by
${\cal M}_{k,k+q}=\delta_{q,0}G_{\rm sp}^{-1}(k)+2\Phi^{\alpha}_{-q}V^{\alpha}_{k,k+q}+\sum_{\sigma=\pm}h^{\mu}_{-q,\sigma}u_{\sigma}^{\mu},$
where $\sigma=\pm$ is the layer index, and $h^{\mu}_{q,\pm}$ is the Fourier transform of the external source $h_{j,\pm}(\tau)$ introduced in Eq. (124) of the main text. The magnetic susceptibility is given by
$\chi^{\mu\nu}_{\alpha}(q)=\frac{\delta^{2}\ln{\cal Z}(h)}{\delta h^{\mu}_{-q,\alpha}\delta h^{\nu}_{q,\alpha}},$ (139)
where the subscript $\alpha=S,A$ denotes the symmetric and antisymmetric combinations between the external source terms from the two layers (see Eq. (125) in the main text). The susceptibility decomposes into two parts, $\chi_{\alpha}(q)=\chi^{{\rm(I)}}_{\alpha}(q)+\chi^{{\rm(II)}}_{\alpha}(q)$, with
$\chi^{{\rm(I)}\mu\nu}_{\alpha}(q)$ $=$ $\frac{1}{2{\cal N}_{D}\beta}\sum_{k}\int{\cal D}[\bar{\Phi},\Phi]\frac{e^{-S_{\rm eff}}}{\cal Z}$
×$\times{\rm tr}\left[{\cal M}^{-1}_{k+q,k}u_{\alpha}^{\mu}{\cal M}^{-1}_{k,k+q}u_{\alpha}^{\nu}\right],$
$\chi^{{\rm(II)}\mu\nu}_{\alpha}(q)$ $=$ $\frac{1}{4{\cal N}_{D}\beta}\sum_{k}\int{\cal D}[\bar{\Phi},\Phi]\frac{e^{-S_{\rm eff}}}{\cal Z}{\rm tr}\left[{\cal M}^{-1}_{k+q,k}u_{\alpha}^{\mu}\right]$ (140)
×$\times{\rm tr}\left[{\cal M}^{-1}_{k,k+q}u_{\alpha}^{\nu}\right]-{\cal N}_{D}\beta m^{\mu}_{\alpha,q}m^{\nu}_{\alpha,-q},$
where
$m^{\mu}_{\alpha,q}\!=\!\frac{-1}{2{\cal N}_{D}\beta}\sum_{k}\!\int\!\!{\cal D}[\bar{\Phi},\Phi]\frac{e^{-S_{\rm eff}}}{\cal Z}{\rm tr}\left[{\cal M}^{-1}_{k+q,k}u_{\alpha}^{\mu}\right]$ (141)
gives rise to the local magnetic moments. The last term in $\chi^{{\rm(II)}}(q)$ eliminates the disconnected diagrams generated from the first term. At this point, one can set $h^{\mu}_{q,\alpha}$ to zero.
The calculation of $\chi^{\mu\nu}_{\alpha}(q)$ can be carried out perturbatively in terms of $1/N$. For this purpose, one needs to expand ${\cal M}^{-1}$, which appears in the integrand of $\chi^{{\rm(I)}\mu\nu}_{\alpha}(q)$ and $\chi^{{\rm(II)}\mu\nu}_{\alpha}(q)$, and $\exp(-S_{\rm eff})$ with respect to the fluctuation field $\Phi$. We have
$\left({\cal M}\right)^{-1}_{k,k+q}$ $=G_{\rm sp}(k)\delta_{q,0}$
$+\sum_{m=1}^{\infty}(-1)^{m}\left(2\Phi^{\alpha}G_{\rm sp}V^{\alpha}\right)^{m}_{k,k+q}G_{\rm sp}(k+q),$
where $\left(2\Phi^{\alpha}G_{\rm sp}V^{\alpha}\right)_{k,k+q}\equiv 2\Phi^{\alpha}_{-q}G_{\rm sp}(k)V^{\alpha}_{k,k+q}$. According to Eq. (LABEL:eq:Seffexp),
$e^{-S_{\rm eff}}=e^{-\frac{4S_{\rm cl}}{N}-\frac{2}{N}\text{Tr}[\ln G_{\rm sp}^{-1}]-S^{(2)}(\bar{\Phi},\Phi)}\sum_{p=0}^{\infty}\frac{(-1)^{p}}{p!}S^{p}_{\rm int},$
where $S_{\rm int}$ is given by Eq. (74). The functional integral over the fluctuation fields is then carried out using Wick’s theorem.
Note that disconnected diagrams generated from this integration are completely canceled out by the perturbative expansion of the partition function ${\cal Z}$. The resulting series expansion can be conveniently represented by Feynman diagrams. The nominal order ${\cal O}(1/N^{L-P})$ of each diagram is determined by the number of internal loops ($L$, defined in Eq. (74)) and the number of RPA propagators.
In Fig. 5, we present the large-$N$ expansion of the magnetic susceptibility up to nominal ${\cal O}(1/N)$. The leading order in Fig. 5 (a), i.e., the SP approximation, corresponds to the leading term in $\chi_{\alpha}^{\rm(I)}$, obtained by taking ${\cal M}^{-1}=G_{\rm sp}$. The $1/N$ diagrams in Fig. 5 (c) correspond to higher-order terms in $\chi_{\alpha}^{\rm(I)}$. The other $1/N$ diagram in Fig. 5 (b) corresponds to the leading term in $\chi_{\alpha}^{\rm(II)}$. As emphasized in the main text, when certain flavors of SBs condense, the $1/N$ diagrams contain singular contributions of ${\cal O}(1)$, which must be considered on an equal footing with the SP approximation.
Figure 14: SP solution of the symmetric DSSF $S^{\mu\mu}_{{\rm sp},S}({\bm{q}},\omega)$ for different values of $g$, where $S^{xx}_{{\rm sp},S}({\bm{q}},\omega)=S^{yy}_{{\rm sp},S}({\bm{q}},\omega)=S^{zz}_{{\rm sp},S}({\bm{q}},\omega)$ is isotropic in QPM phase, and $S^{yy}_{{\rm sp},S}({\bm{q}},\omega)=S^{zz}_{{\rm sp},S}({\bm{q}},\omega)$ refers to the transverse spin components in AFM phase. White lines denote $\varepsilon_{{\rm q},\mu}$ in the QPM phase; extra lines corresponding to $\varepsilon_{{\bm{q}}+{\bm{\pi}},\mu}$ are shown in the Néel AFM phase.
Appendix B Longitudinal DSSF at low frequency
Here we analyze the low frequency behavior of the antisymmetric longitudinal DSSF for the ordering wave vector ${\bm{q}}={\bm{\pi}}$.
In the SP approximation, the longitudinal DSSF in the antisymmetric channel reads
$S^{xx}_{{\rm sp},A}({\bm{\pi}},\omega)=$ $\Theta(\omega)\int\frac{d^{2}{\bm{k}}}{(2\pi)^{2}}(v_{{\bm{k}},0}u_{{\bm{k}},1}+u_{{\bm{k}},0}v_{{\bm{k}},1})^{2}$ (144)
$\delta(\varepsilon_{{\bm{k}},0}+\varepsilon_{{\bm{k}},1}-\omega).$
For small $\omega$, the singlet and triplet boson dispersions are linear near ${\bm{q}}={\bm{0}}$ and ${\bm{q}}={\bm{\pi}}$ with velocities $v_{s}$ and $v_{t}$, respectively. Since the spin operator creates a singlet and a triplet boson with momenta ${\bm{k}}$ and ${\bm{\pi}}-{\bm{k}}$, only $\rvert{\bm{k}}\rvert\sim\omega/(v_{s}+v_{t})$ contributes to the integral. Thus, in the long-wavelength limit we have:
$\varepsilon_{{\bm{k}},0}=v_{s}\rvert{\bm{k}}\rvert(1+{\cal O}(\rvert{\bm{k}}\rvert^{2})),$ (145)
$\varepsilon_{{\bm{k}}+{\bm{\pi}},1}=v_{t}\rvert{\bm{k}}\rvert(1+{\cal O}(\rvert{\bm{k}}\rvert^{2})),$ (146)
$u_{{\bm{k}},0}\approx-v_{{\bm{k}},0}=\sqrt{\frac{\Delta_{s}}{2v_{s}\rvert{\bm{k}}\rvert}}(1+{\cal O}(\rvert{\bm{k}}\rvert^{2})),$ (147)
$u_{{\bm{k}}+{\bm{\pi}},1}\approx-v_{{\bm{k}}+{\bm{\pi}},1}=\sqrt{\frac{\Delta_{t}}{2v_{t}\rvert{\bm{k}}\rvert}}(1+{\cal O}(\rvert{\bm{k}}\rvert^{2})),$ (148)
where $\Delta_{s}=2J^{\prime}{\cal A}$, $\Delta_{t}=2J^{\prime}({\cal A}-{\cal S})$, $v_{s}=\sqrt{\Delta_{s}(\tilde{\lambda}-J)/2}$, $v_{t}=\sqrt{\Delta_{t}\tilde{\lambda}/2}$. For $g=0.8$ considered in the main text, we have $\Delta_{s}\approx 1.0336J$, $\Delta_{t}\approx 2.0542J$, $v_{s}\approx 0.8737J$, and $v_{t}\approx 1.5951J$. By substituting these expressions to Eq. (144), we find
$S^{xx}_{{\rm sp},A}({\bm{\pi}},\omega)={\Theta(\omega)\over 2\pi}{\Delta_{s}\Delta_{t}\over v_{s}v_{t}}{1\over\omega}.$ (149)
The numerical solution presented in the main text matches this form very well in the low-frequency regime. This low frequency behavior is preserved after including the contribution from the counter-diagram shown in Fig. 11 (b). The correction only modifies the prefactor of $1/\omega$.
Appendix C DSSF in the symmetric channel
Figure 15: SP solution of the longitudinal symmetric DSSF $S^{xx}_{{\rm sp},S}({\rm q},\omega)$ for a representative $g=0.8$. White lines denote both $\varepsilon_{{\rm q},\mu}$ and $\varepsilon_{{\bm{q}}+{\bm{\pi}},\mu}$.
Given the exchange symmetry of the two layers (up to a translation in the AFM phase), the symmetric (S) and antisymmetric (A) channels of the magnetic susceptibility, $\chi^{\mu\nu}_{S}(q)$ and $\chi^{\mu\nu}_{A}(q)$, are decoupled. In the main text, we focus on the antisymmetric channel, which exhibits triplon modes in the QPM phase, where their softening leads to the continuous phase transition to the Néel AFM phase, as well as magnon modes and amplitude fluctuation modes in the AFM phase. For completeness, in this Appendix, we present the results for the symmetric channel. Fig. 14 and Fig. 15 show the SP results. Notably, the counter diagram for the symmetric channel does not have singular contributions of ${\cal O}(N^{0})$ and, thus, it is not necessary to correct the leading-order (i.e., the SP) approximation.
The excited states revealed by $\chi^{\mu\nu}_{S}(q)$ are generated by the symmetric spin operator,
$\hat{S}_{jS}^{\mu}=\hat{S}_{j+}^{\mu}+\hat{S}_{j-}^{\mu}=-i\sum_{\nu,\rho=1}^{3}\epsilon^{\mu\nu\rho}\hat{b}_{j,\nu}^{\dagger}\hat{b}_{j,\rho}.$ (150)
This operator is bilinear in two triplet bosons, as it needs to be invariant under the exchange of layers (a bilinear form in singlet bosons is not possible because it would be a scalar under $\mathrm{SU}(2)$ spin rotation). This contrasts with the antisymmetric spin operator, which is bilinear in a singlet and a triplet boson and is therefore odd under the exchange of layers.
In the QPM phase, none of the triplet bosons condense. Thus, each of the two-triplon excitations generated by the symmetric spin operator $\hat{S}_{jS}^{\mu}$ carries spectral weight $\propto 1/{\cal N}_{D}$, forming a continuum of excitations in the DSSF without singular $\delta$ peaks. This results in the DSSF shown in Fig. 14. In contrast, the antisymmetric DSSF shown in Fig. 7 consists of a $\delta$ peak, arising from the creation of a singlet boson in the condensate and a gapped triplet boson, whose spectral weight is of ${\cal O}(1)$ due to the finite condensate fraction of the singlet bosons. It also exhibits a continuum formed by one singlet boson and one triplet boson, which differs from the continuum in the symmetric DSSF.
In the Néel AFM phase, the condensation of the $\mu=1$ triplet boson, $\langle\hat{b}_{j,1}\rangle=\sqrt{n_{c,{\bm{\pi}}}}$, gives rise to local magnetic moments along the $\hat{x}$ direction. For convenience, we denote the longitudinal Schwinger boson (LSB) by ${\hat{b}}_{j,1}$ and the transverse Schwinger bosons (TSBs) by ${\hat{b}}_{j,2/3}$ The transverse symmetric spin operator, $\hat{S}_{jS}^{y/z}$, is bilinear in the LSB and TSB. Since the LSB has a finite condensate fraction, one can replace the LSB operator by $\langle\hat{b}_{j,1}\rangle=\sqrt{n_{c,{\bm{\pi}}}}$, leading to the approximation:
$\hat{S}_{jS}^{y}$ $\simeq i\sqrt{n_{c,{\bm{\pi}}}}(\hat{b}_{j,3}-\hat{b}_{j,3}^{\dagger}),$ (151)
$\hat{S}_{jS}^{z}$ $\simeq i\sqrt{n_{c,{\bm{\pi}}}}(\hat{b}_{j,2}^{\dagger}-\hat{b}_{j,2}).$ (152)
These transverse symmetric spin operators generate quasiparticles (magnons) described by the TSBs, $\hat{b}_{j,2}$ and $\hat{b}_{j,3}$, forming $\delta$ peaks in the symmetric DSSF (see Fig. 14). In the laboratory reference frame, the LSB condenses at the momentum ${\bm{\pi}}$. By momentum conservation, the dispersion of the $\delta$ peak is given by $\varepsilon_{{\bm{\pi}}+{\bm{q}},\mu}$, with $\mu=2,3$, for an external momentum ${\bm{q}}$ in the DSSF, which becomes gapless at the $\Gamma$ point. The spectral weight associated with these $\delta$ peaks is proportional to $n_{c,{\bm{\pi}}}$, which vanishes as the quantum critical point is approached. Fluctuations of the LSB condensate generate a continuum in the symmetric DSSF, arising from a two-particle excitation formed by LSB and TSB, differing again from the antisymmetric channel.
The longitudinal symmetric spin operator, however, is bilinear in two uncondensed TSBs:
$\hat{S}_{jS}^{x}=-i(\hat{b}_{j,2}^{\dagger}\hat{b}_{j,3}-\hat{b}_{j,3}^{\dagger}\hat{b}_{j,2}).$ (153)
It merely excites a two-particle continuum formed by two TSBs, corresponding to a two-magnon continuum since each TSB describes a magnon excitation. This results in the longitudinal DSSF in the symmetric channel shown in Fig. 15. In contrast to the antisymmetric channel shown in Fig. 9, the symmetric channel does not exhibit spurious modes. In other words, the SP description of the longitudinal DSSF in the symmetric channel is qualitatively correct.
As we already mentioned, the counter diagram shown in Fig. 11(b) vanishes for the symmetric channel. Specifically, the cross susceptibility between the external symmetric spin operators and the auxiliary field operators must vanish. For the QPM phase and the transverse spin components in the Néel AFM phase, the same symmetry argument provided in the main text for the antisymmetric channel still applies. For the longitudinal spin components in the Néel AFM phase, while both the longitudinal spin operator $\hat{S}_{jS}^{x}$ and the auxiliary field operators are invariant under the residual $\mathrm{U}(1)$ spin rotation, their cross susceptibility still vanishes because $\hat{S}_{jS}^{x}$ creates two uncondensed TSBs that preserve all symmetries of the model. Note that, for the cross susceptibility to be finite, the intermediate two-particle state must break certain symmetries of the system. As in the antisymmetric channel, the longitudinal spin operator $\hat{S}_{jA}^{x}$ creates a singlet boson and an LSB, both of which condense in a hybridized energy level that results in the formation of local magnetic moments.